### v1

# Descending finite projective modules from a Novikov ring

Abstract.
We prove a descent result for finite projective modules, motivated by a question in perfectoid geometry. Given a commutative ring $A$, we formulate a descent problem for descending a finite projective module over the Novikov ring with coefficients in $A$ to a finite projective module over $A$. The main theorem of this paper is that all such descent data are effective. As an application, we prove for every perfect $\mathbb{F}_{p}$-algebra $A$, a vector bundle on $\operatorname{Spd}A$ always descends to a vector bundle on $\operatorname{Spec}A$.
1. Introduction
The goal of this paper is to prove the following descent result. We refer the reader to Section 2 for the definition of the rings and maps involved.
Theorem 1.1 (Theorem 4.9).
Let $A$ be a commutative ring and let $\Gamma\subseteq\mathbb{R}$ be an additive submonoid. Then the category of descent data of finite projective modules along the semi-cosimplicial diagram
$A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\rightrightarrows\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!)$
is equivalent to the category of finite projective $A$-modules.
Remark 1.2.
In general, the ring $A(\!(t^{\Gamma},u^{\Gamma})\!)$ is much larger than the tensor product $A(\!(t^{\Gamma})\!)\otimes_{A}A(\!(t^{\Gamma})\!)$, which is usually where the isomorphism of the descent datum is defined. As a consequence, standard descent results for finite projective modules are inapplicable to this setting.
The theorem is motivated by question in perfectoid geometry, namely that of descending a vector bundle from $\operatorname{Spd}A$ to a vector bundle on $\operatorname{Spec}A$, where $A$ is a discrete perfect ring in characteristic $p$. Using Theorem 4.9 with $\Gamma=\mathbb{Z}[p^{-1}]$, we deduce the following result as an immediate consequence.
Theorem 1.3 (Corollary 5.2, 5.4).
Let $A$ be a perfect $\mathbb{F}_{p}$-algebra. Then the natural functor
{vector bundles on $\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
is an exact tensor equivalence. Similarly, the natural functor
[…]
is an equivalence.
Remark 1.4.
It is not true in general that a v-sheaf étale over $\operatorname{Spd}A$ descends to a scheme étale over $\operatorname{Spec}A$. For example, consider the perfect $\mathbb{F}_{p}$-algebra $A=\mathbb{F}_{p}[t^{1/p^{\infty}}]$. The associated v-sheaf is $(\operatorname{Spd}A)(R,R^{+})=R^{+}$, and admits a v-subsheaf $X(R,R^{+})=R^{\circ\circ}$. Then $X\to\operatorname{Spd}A$ is an open embedding, hence étale, but does not correspond to an étale $A$-scheme. Indeed, for every closed point $x\in\operatorname{Spec}A$, the fiber $\operatorname{Spd}k(x)\times_{\operatorname{Spd}A}X$ is empty when $x\neq 0$ and is $\operatorname{Spd}\mathbb{F}_{p}$ when $x=0$.
Remark 1.5.
The first part of the theorem is false when we replace $\operatorname{Spec}A$ with an affinoid rigid analytic variety over a perfectoid field of characteristic zero, by the main result of [Heu22]. On the other hand, the second part of the theorem remains true by [Sch22, Lemma 15.6].
The second part of the theorem, Corollary 5.4, will be used in a future work [DKvHZ24] to construct integral models of Hodge type Shimura varieties at parahoric level. Specifically, it allows one to construct a finite étale cover of the special fiber a Shimura variety from the finite étale cover of the associated v-sheaf.
Remark 1.6.
Ian Gleason announced via private communication an alternative proof of Corollary 5.4, which works by comparing the category of étale $\mathbb{F}_{\ell}$-local systems on $\operatorname{Spd}A$ and $\operatorname{Spec}A$.
1.1. Outline of proof
We sketch the main ideas that go into the proof of Theorem 4.9. In Section 3, we introduce the notion of a Novikov isocrystal, which somewhat easier to control than a descent datum as it involves only one variable. In Subsection 4.1, we construct a functor from the category of Novikov descent data for $\Gamma=\mathbb{R}$ to the category of Novikov isocrystals, and prove that this functor is always fully faithful. This allows us to reduce the effectivity of a Novikov descent datum into a similar statement about the associated Novikov isocrystal.
In Subsection 4.2, we prove effectivity of Novikov descent data when $A$ is reduced and $\Gamma=\mathbb{R}$. We first take care the case when $A$ is an algebraically closed field by using the structure theory of Novikov isocrystals developed in Section 3. Next, we check that when $A$ is an arbitrary product of algebraically closed fields, effectivity over $A$ follows from that over each field. This allows to deduce effectivity for a general reduced ring $A$ by embedding $A$ into the product of the algebraic closures of its residue fields.
In Subsection 4.3, we remove the hypothesis that $A$ is reduced. We first check that effectivity of a Novikov descent datum can be extended along a finite-order nilpotent thickening. On the other hand, it turns out that we can reduce an arbitrary nilpotent thickening to a finite-order thickening by studying the associated Novikov isocrystal. Combining the two allows us to prove effectivity in the case of a general ring $A$ with $\Gamma=\mathbb{R}$. Finally, in Subsection 4.4, we remove the hypothesis that $\Gamma=\mathbb{R}$ and prove Theorem 4.9 in its full generality.
1.2. Acknowledgements
I would like to thank Patrick Daniels, Pol van Hoften, and Mingjia Zhang for help formulating the problem, which arose in our work on integral models of Shimura varieties, as well as insightful conversations. I would also like to thank my advisor Richard Taylor for his encouragements and diligence during our meetings, where we carefully went through parts of the argument. I also would like to thank Brian Conrad, Ian Gleason, and Kiran Kedlaya for helpful comments and suggestions.
2. Novikov rings
Definition 2.1.
Let $\Gamma\subseteq\mathbb{R}$ be an additive submonoid, and let $A$ be a commutative ring. We define the Novikov ring with exponents in $\Gamma$ and coefficients in $A$ to be the $A$-algebra
$A(\!(t^{\Gamma})\!)=\biggl{\{}\sum_{i\in I}a_{i}t^{d_{i}}:a_{i}\in A,d_{i}\in\Gamma,\{i\in I:d_{i}<C\}\text{ is finite for all }C\in\mathbb{R}\biggr{\}}.$
For $M$ an $A$-module, we similarly define the $A(\!(t^{\Gamma})\!)$-module
$M(\!(t^{\Gamma})\!)=\biggl{\{}\sum_{i\in I}m_{i}t^{d_{i}}:m_{i}\in M,d_{i}\in\Gamma,\{i\in I:d_{i}<C\}\text{ is finite for all }C\in\mathbb{R}\biggr{\}}.$
Remark 2.2.
Such rings naturally appear as coefficient rings in Lagrangian Floer theory [FOOO09], and are named after Novikov for his use of a formal variable in the coefficient ring of a homology theory to keep track of topological invariants [Nov81].
We note that the functor $M\mapsto M(\!(t^{\Gamma})\!)$ sending $A$-modules to $A(\!(t^{\Gamma})\!)$-modules commutes with limits and finite colimits. In particular, when $M$ is a finitely presented $A$-module, the natural map
$M\otimes_{A}A(\!(t^{\Gamma})\!)\to M(\!(t^{\Gamma})\!)$
is an isomorphism. As a consequence, for any other $A$-module $N$ there is a canonical isomorphism
$\operatorname{Hom}_{A(\!(t^{\Gamma})\!)}(M(\!(t^{\Gamma})\!),N(\!(t^{\Gamma})\!))\cong\operatorname{Hom}_{A}(M,N)(\!(t^{\Gamma})\!)$
of $A(\!(t^{\Gamma})\!)$-modules.
To formulate our descent problem, we define a version of the Novikov ring with multiple variables.
Definition 2.3.
Let $\Gamma\subseteq\mathbb{R}$ be an additive submonoid, let $A$ be a commutative ring, and let $M$ be an $A$-module. We define the generalized Novikov ring with $n$ variables
$A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)=\left\{\sum_{i\in I}a_{i}t_{1}^{d_{i,1}}\dotsm t_{n}^{d_{i,n}}:\begin{matrix}a_{i}\in A,d_{i,j}\in\Gamma,\{i\in I:\sum_{j}s_{j}d_{i,j}<C\}\text{ is}\\ \text{finite for all }C\in\mathbb{R}\text{ and }s_{1},\dotsc,s_{n}\in\mathbb{R}_{>0}\end{matrix}\right\}$
and the $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$-module
$M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)=\left\{\sum_{i\in I}m_{i}t_{1}^{d_{i,1}}\dotsm t_{n}^{d_{i,n}}:\begin{matrix}m_{i}\in M,d_{i,j}\in\Gamma,\{i\in I:\sum_{j}s_{j}d_{i,j}<C\}\text{ is}\\ \text{finite for all }C\in\mathbb{R}\text{ and }s_{1},\dotsc,s_{n}\in\mathbb{R}_{>0}\end{matrix}\right\},$
where we note that the definitions of $A(\!(t^{\Gamma})\!)$ and $M(\!(t^{\Gamma})\!)$ agree with that of Definition 2.1.
We define $A$-algebra homomorphisms
$\pi_{1}^{\ast}$ $\colon A(\!(t^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma})\!);$ $t$ t,maps-$\mapsto t,$
$\pi_{2}^{\ast}$ $\colon A(\!(t^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma})\!);$ $t$ $\mapsto u,$
$\pi_{12}^{\ast}$ $\colon A(\!(t^{\Gamma},u^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!);$ $t$ $\mapsto t,u\mapsto u,$
$\pi_{13}^{\ast}$ $\colon A(\!(t^{\Gamma},u^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!);$ $t$ $\mapsto t,u\mapsto v,$
$\pi_{23}^{\ast}$ $\colon A(\!(t^{\Gamma},u^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!);$ $t$ $\mapsto u,u\mapsto v.$
They form an augmented semi-cosimplicial diagram of $A$-algebras
$A\to A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\rightrightarrows\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!).$
We will also use the notation $\rho_{1}^{\ast}=\pi_{12}^{\ast}\circ\pi_{1}^{\ast}=\pi_{13}^{\ast}\circ\pi_{1}^{\ast}$ and similarly $\rho_{2}^{\ast}=\pi_{12}^{\ast}\circ\pi_{2}^{\ast}=\pi_{23}^{\ast}\circ\pi_{1}^{\ast}$ and $\rho_{3}^{\ast}=\pi_{13}^{\ast}\circ\pi_{2}^{\ast}=\pi_{23}^{\ast}\circ\pi_{2}^{\ast}$. For a ring homomorphism $f^{\ast}\colon A\to B$ and an $A$-module $M$, we will also denote $f^{\ast}M=M\otimes_{A,f^{\ast}}B$, and abuse notation to write $f^{\ast}\colon M\to f^{\ast}M$ for the map $m\mapsto m\otimes 1$.
Definition 2.4.
A Novikov descent datum for a ring $A$ is a pair $(M,\varphi)$ where $M$ is a finite projective module over $A(\!(t^{\Gamma})\!)$ and
$\varphi\colon\pi_{1}^{\ast}M\xrightarrow{\cong}\pi_{2}^{\ast}M$
is a $A(\!(t^{\Gamma},u^{\Gamma})\!)$-linear automorphism satisfying the cocycle condition
[…]
We denote by $\mathsf{NovDD}(A,\Gamma)$ the category of Novikov descent data.
Denote by $\mathsf{Vect}(A)$ the category of finite projective $A$-modules. There is a natural functor
$\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma);\quad M_{0}\mapsto(M_{0}\otimes_{A}A(\!(t^{\Gamma})\!),\operatorname{id}_{M_{0}\otimes_{A}A(\!(t^{\Gamma},u^{\Gamma})\!)}),$
which will turn out to be an equivalence by Theorem 4.9.
Definition 2.5.
For an element $x\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ expressed as
$x=\sum_{i\in I}a_{i}t_{1}^{d_{i,1}}\dotsm t_{n}^{d_{i,n}}$
where $a_{i}\neq 0$ for all $i\in I$ and $(d_{i,1},\dotsc,d_{i,n})\in\Gamma^{n}$ are all distinct, we define the support of $x$ as
$\operatorname{supp}(x)=\{(d_{i,1},\dotsc,d_{i,n}):i\in I\}\subseteq\Gamma^{n}.$
We similarly define the support of an element of $M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ for $M$ an $A$-module.
The support of $0\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ is empty. We always have $\operatorname{supp}(x+y)\subseteq\operatorname{supp}(x)\cup\operatorname{supp}(y)$ and $\operatorname{supp}(xy)\subseteq\operatorname{supp}(x)+\operatorname{supp}(y)$, where $+$ is the Minkowski sum defined as $X+Y=\{x+y:x\in X,y\in Y\}$. By definition, for every element $x\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ or $x\in M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$, the intersection of $\operatorname{supp}(x)$ with the half-space
$\{(d_{1},\dotsc,d_{n}):s_{1}d_{1}+\dotsb+s_{n}d_{n}<C\}\subseteq\mathbb{R}^{n}$
is finite whenever $s_{1},\dotsc,s_{n}>0$.
Proposition 2.6.
The functor $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma)$ is fully faithful.
Proof.
Given finite projective $A$-modules $M_{0}$ and $N_{0}$, let
$f\colon M_{0}(\!(t^{\Gamma})\!)\to N_{0}(\!(t^{\Gamma})\!)$
be a $A(\!(t^{\Gamma})\!)$-linear map compatible with the descent data on both sides. Viewing $f$ as an element of $\operatorname{Hom}_{A}(M_{0},N_{0})(\!(t^{\Gamma})\!)$, we see that compatibility with descent data translates to
$\pi_{1}^{\ast}f=\pi_{2}^{\ast}f\in\operatorname{Hom}_{A}(M_{0},N_{0})(\!(t^{\Gamma},u^{\Gamma})\!).$
We see that $\operatorname{supp}(\pi_{1}^{\ast}f)=\operatorname{supp}(f)\times\{0\}$ while $\operatorname{supp}(\pi_{2}^{\ast}f)=\{0\}\times\operatorname{supp}(f)$. This immediately shows that $\operatorname{supp}(f)\subseteq\{0\}$ and hence $f\in\operatorname{Hom}_{A}(M_{0},N_{0})$. ∎
2.1. Topological rings and modules
We start by reviewing general facts about finite projective modules over topological rings, which we do not assume are Hausdorff. (In practice, we will apply the results to $A(\!(t^{\mathbb{R}})\!)$, which is always Hausdorff.) A topological ring is a ring $R$ together with a topology such that the addition and multiplication maps $R\times R\to R$ are continuous, where $R\times R$ is given the product topology. A topological module over a topological ring $R$ is an $R$-module $M$ together with a topology such that the addition map $M\times M\to M$ and the scalar multiplication map $R\times M\to M$ are continuous.
Definition 2.7.
Let $R$ be a topological ring, and let $M$ be a finite projective $R$-module. The canonical topology on $M$ is defined as the coarsest topology under which all $R$-linear homomorphisms $M\to R$ are continuous.
Lemma 2.8.
Let $R$ be a topological ring, and let $M,N$ be finite projective $R$-modules.
  1. (1)
    The canonical topology on $R$ as an $R$-module agrees with the topology on $R$.
  2. (2)
    The canonical topology on $M$ makes $M$ into a topological $R$-module.
  3. (3)
    The canonical topology on $M\oplus N$ is the product topology of the canonical topologies on $M$ and $N$.
  4. (4)
    Every $R$-linear map $M\to N$ is continuous with respect to the canonical topologies on $M$ and $N$.
Proof.
The statement (4) follows from unraveling the definition of the coarsest topology; the canonical topology on $N$ is generated by preimages of open sets $U\subseteq R$ along $R$-linear maps $N\to R$, and its preimage in $M$ is the preimage of $U$ under the composition $M\to N\to R$, hence open. For (1), we note that the canonical topology is at least as fine as the ring topology, since $\operatorname{id}\colon R\to R$ is linear. On the other hand, it is fine enough because every scaling $r\colon R\to R$ for $r\in R$ is continuous. For (3), we similarly note that the canonical topology is at least as fine as a product topology, because the projection maps $M\oplus N\to M$ and $M\oplus N\to N$ are continuous. It is fine enough, because any $M\oplus N\to R$ can be written as a composition
$M\times N\xrightarrow{f\times g}R\times R\xrightarrow{+}R$
where $f\colon M\to R$ and $g\colon N\to R$ are $R$-linear, hence continuous. Finally, for (2), we choose a splitting $M\oplus M^{\prime}\cong R^{\oplus n}$ of $R$-modules. By (4), both the inclusion $M\hookrightarrow R^{\oplus n}$ and the projection $R^{\oplus n}\twoheadrightarrow M$ are continuous. The addition and scalar multiplication maps on $M$ can be written as the composition
$M\times M\hookrightarrow R^{\oplus n}\times R^{\oplus n}\xrightarrow{+}R^{\oplus n}\twoheadrightarrow M,\quad R\times M\hookrightarrow R\times R^{\oplus n}\xrightarrow{\times}R^{\oplus n}\twoheadrightarrow M$
of continuous maps. ∎
We focus on the case when there is one variable and the exponent group is $\Gamma=\mathbb{R}$. For convenience, for $A$ a commutative ring and $M$ an $A$-module, we denote
$A[[t^{\mathbb{R}}]]=A(\!(t^{\mathbb{R}_{\geq 0}})\!),\quad M[[t^{\mathbb{R}}]]=M(\!(t^{\mathbb{R}_{\geq 0}})\!).$
Definition 2.9.
We define the $t$-adic topology on $A(\!(t^{\mathbb{R}})\!)$ to be the topology generated by $x+t^{d}A[[t^{\mathbb{R}}]]$ for varying $x\in A(\!(t^{\mathbb{R}})\!)$ and $d\in\mathbb{R}$. We similarly define the $t$-adic topology on $M(\!(t^{\mathbb{R}})\!)$ to be the topology generated by $x+t^{d}M[[t^{\mathbb{R}}]]$ for varying $x\in M(\!(t^{\mathbb{R}})\!)$ and $d\in\mathbb{R}$.
It can be readily verified that $A(\!(t^{\mathbb{R}})\!)$ is a topological ring. Indeed, it is a complete Tate ring in the sense of [Hub93], where $A[[t^{\mathbb{R}}]]$ is a ring of definition. It can also be checked that $M(\!(t^{\mathbb{R}})\!)$ is a complete topological $A(\!(t^{\mathbb{R}})\!)$-module.
If $M$ is a finite projective $A$-module, then $M(\!(t^{\mathbb{R}})\!)$ is a finite projective $A(\!(t^{\mathbb{R}})\!)$-module, and moreover the $t$-adic topology on $M(\!(t^{\mathbb{R}})\!)$ agrees with the canonical topology. We now a prove a converse statement.
Lemma 2.10.
Let $A$ be a commutative ring, and let $M$ be an $A$-module. Assume that $M(\!(t^{\mathbb{R}})\!)$ is a finite projective $A(\!(t^{\mathbb{R}})\!)$-module, and moreover the $t$-adic topology on $M(\!(t^{\mathbb{R}})\!)$ agrees with the canonical topology. Then $M$ is a finite projective $A$-module.
Proof.
We first show that $M$ is finitely generated as an $A$-module. We choose a finite collection of generators $x_{1},\dotsc,x_{n}\in M(\!(t^{\mathbb{R}})\!)$ as an $A(\!(t^{\mathbb{R}})\!)$-module. Since $M(\!(t^{\mathbb{R}})\!)$ is projective, the natural map
$A(\!(t^{\mathbb{R}})\!)^{\oplus n}\xrightarrow{(x_{1},\dotsc,x_{n})}M(\!(t^{\mathbb{R}})\!)$
splits and hence is open by Lemma 2.8. This implies that there exists a large enough $N$ for which
$t^{N}M(\!(t^{\mathbb{R}})\!)\subseteq\sum_{i=1}^{n}x_{i}A[[t^{\mathbb{R}}]].$
In particular, for every $m\in M$, the element $t^{N}m\in M(\!(t^{\mathbb{R}})\!)$ can be written as $A[[t^{\mathbb{R}}]]$-linear combination of $x_{1},\dotsc,x_{n}$. Therefore once we write $x_{i}=\sum_{j}m_{ij}t^{d_{ij}}$, the collection of $m_{ij}$ with $d_{ij}\leq N$ generate $M$, and so $M$ is finitely generated.
We now choose an $A$-linear surjection
$0\to K\to A^{\oplus m}\to M\to 0$
It follows that
$0\to K(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}})\!)^{\oplus m}\to M(\!(t^{\mathbb{R}})\!)\to 0$
is short exact, and hence split. This implies that $K(\!(t^{\mathbb{R}})\!)$ is finite projective over $A(\!(t^{\mathbb{R}})\!)$. Moreover the canonical topology of $K(\!(t^{\mathbb{R}})\!)$ is the subspace topology induced from $A(\!(t^{\mathbb{R}})\!)^{\oplus m}$, again by Lemma 2.8, and hence agrees with the $t$-adic topology. By applying the same argument, we conclude that $K$ is also finitely generated, and therefore $M$ is finitely presented.
This shows that the natural map
$M\otimes_{A}A(\!(t^{\mathbb{R}})\!)\to M(\!(t^{\mathbb{R}})\!)$
is an isomorphism. Since $A\hookrightarrow A(\!(t^{\mathbb{R}})\!)$ splits as a map of $A$-modules, it is universally injective. Then [Sta24, Theorem 08XD] applies to show that $M$ is a finite projective $A$-module. ∎
3. Novikov isocrystals
We fix a real number $\lambda>1$. For each commutative ring $A$, we define the $A$-algebra automorphism
$F\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}})\!);\quad\sum_{i\in I}a_{i}t^{d_{i}}\mapsto\sum_{i\in I}a_{i}t^{\lambda d_{i}}.$
Definition 3.1.
A Novikov isocrystal over a ring $A$ is a pair $(M,F_{M})$, where $M$ is a finite projective $A(\!(t^{\mathbb{R}})\!)$-module and $F_{M}\colon M\to M$ is a $F$-semilinear bijection. A morphism $f\colon(M,F_{M})\to(N,F_{N})$ between two Novikov isocrystals is a $A(\!(t^{\mathbb{R}})\!)$-linear map $f\colon M\to N$ satisfying $f\circ F_{M}=F_{N}\circ f$. We denote by $\mathsf{NovIsoc}(A)$ the category of Novikov isocrystals over $A$.
Remark 3.2.
The name “Novikov isocrystal” comes from the analogy with the notion of a F-isocrystal [Man63]. For $k$ a perfect field in characteristic $p>0$, a F-isocrystal over $k$ is a $W(k)[p^{-1}]$-vector space $M$ together with a $F$-semilinear bijection $F_{M}\colon M\to M$, where $W(k)$ is the ring of $p$-typical Witt vectors on $k$ and $F\colon W(k)[p^{-1}]\to W(k)[p^{-1}]$ is the canonical lift of the absolute Frobenius.
Note that a ring homomorphism $A\to B$ induces a $F$-equivariant ring homomorphism $A(\!(t^{\mathbb{R}})\!)\to B(\!(t^{\mathbb{R}})\!)$, and hence base changing along this map defines a functor $\mathsf{NovIsoc}(A)\to\mathsf{NovIsoc}(B)$. There is also a natural functor
$\mathsf{Vect}(A)\to\mathsf{NovIsoc}(A);\quad M_{0}\mapsto(M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!),\operatorname{id}_{M_{0}}\otimes F).$
Proposition 3.3.
For every ring $A$, the functor $\mathsf{Vect}(A)\to\mathsf{NovIsoc}(A)$ is fully faithful.
Proof.
Let $M_{0}$ and $N_{0}$ be two finite projective $A$-modules, and let $f\colon M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)\to N_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)$ be a $A(\!(t^{\mathbb{R}})\!)$-linear map satisfying $f\circ(\operatorname{id}_{M_{0}}\otimes F)=(\operatorname{id}_{N_{0}}\otimes F)\circ f$. Regarding $f$ as an element of
$\operatorname{Hom}_{A}(M_{0},N_{0})\otimes_{A}A(\!(t^{\mathbb{R}})\!)=\operatorname{Hom}_{A}(M_{0},N_{0})(\!(t^{\mathbb{R}})\!),$
we see that $f$ must be fixed under the automorphism $\operatorname{id}_{\operatorname{Hom}_{A}(M_{0},N_{0})}\otimes F$, and hence $\operatorname{supp}(f)=\lambda\operatorname{supp}(f)$. On the other hand, the intersection $[-1,1]\cap\operatorname{supp}(f)$ is finite, and therefore $\operatorname{supp}(f)\subseteq\{0\}$. This shows that $f\in\operatorname{Hom}_{A}(M_{0},N_{0})$. ∎
Remark 3.4.
The functor $\mathsf{Vect}(A)\to\mathsf{NovIsoc}(A)$ is far from being essentially surjective. If we denote by $\mathsf{VectAut}(A)$ the category of pairs $(M_{0},F_{M_{0}})$ where $M_{0}$ is a finite projective $A$-module and $F_{M_{0}}\colon M_{0}\to M_{0}$ is an $A$-linear automorphism, there is a fully faithful functor $\mathsf{VectAut}(A)\to\mathsf{NovIsoc}(A)$ sending $(M_{0},F_{M_{0}})$ to $(M_{0}(\!(t^{\mathbb{R}})\!),F_{M_{0}}\otimes F)$. Even then, the functor $\mathsf{VectAut}(A)\to\mathsf{NovIsoc}(A)$ is not necessarily essentially surjective. Already for $k$ a field, the Novikov isocrystal
$M=k(\!(t^{\mathbb{R}})\!)^{\oplus 2},\quad F_{M}=\begin{pmatrix}1&t^{-1}\\ 0&1\end{pmatrix}\circ(F\oplus F)$
is not contained in the essential image of $\mathsf{VectAut}(k)$.
3.1. Some results on isocrystals
In this subsection, we provide some results that allow us to control the structure of a Novikov isocrystal. These results will later be used in the proof of Theorem 4.9.
Proposition 3.5.
Let $k$ be an algebraically closed field. Then objects of the form $(k(\!(t^{\mathbb{R}})\!),cF)$ for some $c\in k^{\times}$ generate $\mathsf{NovIsoc}(k)$ under extensions.
Proof.
We imitate the argument in the proof of the Dieudonné–Manin classification [Man63]. Because $k(\!(t^{\mathbb{R}})\!)$ is a field, it suffices to prove that for every nonzero Novikov isocrystal $(M,F_{M})$, there exists a nonzero element $m\in M$ and $c\in k^{\times}$ satisfying $F_{M}(m)=cm$. Fix an arbitrary nonzero $m\in M$ and find $n\geq 1$ such that $m,F_{M}(m),\dotsc,F_{M}^{n-1}(m)$ are linearly independent while
$F_{M}^{n}(m)=a_{0}m+a_{1}F_{M}(m)+\dotsb+a_{n-1}F_{M}^{n-1}(m)$
for $a_{1},\dotsc,a_{n}\in k(\!(t^{\mathbb{R}})\!)$. Replacing $m$ with $t^{r}m$ for a real number $r\in\mathbb{R}$ has the effect of replacing $a_{i}$ with $t^{r(\lambda^{n}-\lambda^{i})}$. Thus by choosing $r$ appropriately, we may further assume that $a_{1},\dotsc,a_{n}\in k[[t^{\mathbb{R}}]]$ with $a_{i}\in k[[t^{\mathbb{R}}]]^{\times}$ for some $10\leq i\leq n-$ 1.
Our goal is to find elements $b_{0},\dotsc,b_{n-1}\in k(\!(t^{\mathbb{R}})\!)$, not all zero, and a scalar $c\in k^{\times}$ satisfying the equation
$F_{M}(b_{0}m+b_{1}F_{M}(m)+\dotsb+b_{n-1}F_{M}^{n-1}(m))=cb_{0}m+cb_{1}F_{M}(m)+\dotsb+cb_{n-1}F_{M}^{n-1}(m).$
This amounts to the system of equations
$F(b_{n-1})a_{0}=cb_{0},\;F(b_{0})+F(b_{n-1})a_{1}=cb_{1},\;\dotsc,\;F(b_{n-2})+F(b_{n-1})a_{n-1}=cb_{n-1},$
which can be rewritten as a single equation
$F^{n}(b)F^{n-1}(a_{0})+cF^{n-1}(b)F^{n-2}(a_{1})+\dotsb+c^{n-1}F(b)a_{n-1}=c^{n}b$
in terms of $b=b_{n-1}$.
We claim that this equation has a nonzero solution in $k[[t^{\mathbb{R}}]]$. Let $\bar{a}_{i}\in k$ be the constant terms of $a_{i}\in k[[t^{\mathbb{R}}]]$. Since we are assuming that $\bar{a}_{i}\neq 0$ for some $11\leq i\leq n-$ 1 and that $k$ is algebraically closed, there exists a $c\in k^{\times}$ for which
$\bar{a}_{0}+c\bar{a}_{1}+\dotsb+c^{n-1}\bar{a}_{n-1}=c^{n}.$
Write
$g(b)=c^{-n}F^{n}(b)F^{n-1}(a_{0})+c^{-n+1}F^{n-1}(b)F^{n-2}(a_{1})+\dotsb+c^{-1}F(b)$
so that we need to find $b\in k[[t^{\mathbb{R}}]]$ with $g(b)=b$. Set $b^{(0)}=1$ and $b^{(j+1)}=g(b^{(j)})$. By construction of $c$, the constant term of $b^{(1)}=g(b^{(0)})$ is $1$, and hence there exists a positive real number $\epsilon>0$ for which $b^{(1)}-b^{(0)}\in t^{\epsilon}k[[t^{\mathbb{R}}]]$. On the other hand, we observe that $x-y\in t^{d}k[[t^{\mathbb{R}}]]$ implies $g(x)-g(y)\in t^{\lambda d}k[[t^{\mathbb{R}}]]$. This shows that $b^{(j+1)}-b^{(j)}\in t^{\lambda^{i}\epsilon}k[[t^{\mathbb{R}}]]$ and therefore
$b=\lim_{i\to\infty}b^{(j)}\in 1+t^{\epsilon}k[[t^{\mathbb{R}}]]$
exists and satisfies $g(b)=b$. ∎
Lemma 3.6.
Let $A$ be a commutative ring, and let $M_{0}$ be an $A$-module. Consider the topological $A(\!(t^{\mathbb{R}})\!)$-module $M_{0}(\!(t^{\mathbb{R}})\!)$, where $M_{0}(\!(t^{\mathbb{R}})\!)$ has the $t$-adic topology. If $S\subseteq M_{0}(\!(t^{\mathbb{R}})\!)$ is a closed $A$-submodule that is stable under the map
$F\colon M_{0}(\!(t^{\mathbb{R}})\!)\to M_{0}(\!(t^{\mathbb{R}})\!);\quad\sum_{i}m_{i}t^{d_{i}}\mapsto\sum_{i}m_{i}t^{\lambda d_{i}},$
and also stable under multiplication by $t^{d}$ for every $d\in\mathbb{R}$, then there exists an $A$-submodule $S_{0}\subseteq M_{0}$ for which $S=S_{0}(\!(t^{\mathbb{R}})\!)$.
Proof.
Choose any element $s\in S$ and write
$s=\sum_{i\geq 1}m_{i}t^{d_{i}}\in M_{0}(\!(t^{\mathbb{R}})\!),\quad m_{i}\in M_{0},\quad d_{1}<d_{2}<\dotsb.$
Since $S$ is closed and stable under $F$, we have
$m_{1}=\lim_{k\to\infty}F^{k}(t^{-d_{1}}s)\in S.$
Then $\sum_{i\geq 2}m_{i}t^{d_{i}}=s-m_{1}t^{d_{1}}\in S$ and similarly $m_{2}\in S$. Iterating the process shows that $m_{i}\in S$ for all $i$, and therefore all coefficients that appear in the power series expansion of elements of $S$ are in $S$. Let $S_{0}\subseteq M_{0}$ be the subset of all such coefficients that appear in elements of $S$, so that $S\subseteq S_{0}(\!(t^{\mathbb{R}})\!)$. The above analysis shows that $S_{0}\subseteq S$, and by writing each element of $S_{0}(\!(t^{\mathbb{R}})\!)$ as a convergent sum, we moreover obtain $S=S_{0}(\!(t^{\mathbb{R}})\!)$. ∎
Proposition 3.7.
Let $A\hookrightarrow B$ be an injective ring homomorphism, and let $(M,F_{M})\in\mathsf{NovIsoc}(A)$ be a Novikov isocrystal with the property that its base change to $B$ lies in the essential image of $\mathsf{Vect}(B)\to\mathsf{NovIsoc}(B)$. Then $(M,F_{M})$ lies in the essential image of $\mathsf{Vect}(A)\to\mathsf{NovIsoc}(A)$.
Proof.
By assumption, there exists a finite projective $B$-module $N_{0}$ together with a $B(\!(t^{\mathbb{R}})\!)$-linear isomorphism
$M\otimes_{A(\!(t^{\mathbb{R}})\!)}B(\!(t^{\mathbb{R}})\!)\cong N_{0}\otimes_{B}B(\!(t^{\mathbb{R}})\!)$
such that the action of $F_{M}\otimes F$ on the left hand side agrees with the action of $\operatorname{id}_{N_{0}}\otimes F$ on the right hand side.
Since $M$ is flat over $A(\!(t^{\mathbb{R}})\!)$ and $A(\!(t^{\mathbb{R}})\!)\to B(\!(t^{\mathbb{R}})\!)$ is injective, we may use the composition
$M\hookrightarrow M\otimes_{A(\!(t^{\mathbb{R}})\!)}B(\!(t^{\mathbb{R}})\!)\cong N_{0}\otimes_{B}B(\!(t^{\mathbb{R}})\!)=N_{0}(\!(t^{\mathbb{R}})\!)$
to regard $M$ as a $A(\!(t^{\mathbb{R}})\!)$-submodule of $N_{0}(\!(t^{\mathbb{R}})\!)$ that is stable under the action of $\operatorname{id}_{N_{0}}\otimes F$. Note that $A(\!(t^{\mathbb{R}})\!)\hookrightarrow B(\!(t^{\mathbb{R}})\!)$ is a closed subring with the subspace topology. By writing $M$ as a direct summand of a finite free module, we see that $M$ is also closed in $N_{0}(\!(t^{\mathbb{R}})\!)$, and moreover the canonical topology on $M$ agrees with the subspace topology from $N_{0}(\!(t^{\mathbb{R}})\!)$.
Using Lemma 3.6, we see that there exists an $A$-submodule $M_{0}\subseteq N_{0}$ for which $M=M_{0}(\!(t^{\mathbb{R}})\!)$. On the other hand, we observe from above that the canonical topology on $M$ agrees with the $t$-adic topology. It therefore follows from Lemma 2.10 that $M_{0}$ is a finite projective $A$-module. ∎
Proposition 3.8.
Let $A$ be a commutative ring, and let $(M,F_{M})\in\mathsf{NovIsoc}(A)$ be a Novikov isocrystal. Consider a finitely generated $A[[t^{\mathbb{R}}]]$-submodule $M^{\circ}\subseteq M$ with the property that $M^{\circ}[t^{-1}]=M$. Assume that for each $m^{\circ}\in M^{\circ}$ there exists a real number $\epsilon>0$ for which
$F_{M}(m^{\circ})-m^{\circ}\in t^{\epsilon}M^{\circ}.$
Then $(M,F_{M})$ is in the essential image of $\mathsf{Vect}(A)\to\mathsf{NovIsoc}(A)$.
Proof.
Consider the canonical topology on $M$. We first claim that $t^{d}M^{\circ}$ for $d\in\mathbb{R}$ form a neighborhood basis of $0\in M$. To see this, we choose the finite set of generators of $M^{\circ}$ so that
$f\colon A(\!(t^{\mathbb{R}})\!)^{\oplus n}\twoheadrightarrow M$
is surjective with the image of $A[[t^{\mathbb{R}}]]^{\oplus n}$ being $M^{\circ}$. Since $M$ is projective, it split and hence $f$ is open by Lemma 2.8. This immediately implies that $M^{\circ}$ is open, and hence $t^{d}M^{\circ}$ are also open. On the other hand, for every open neighborhood $0\in U$ in $M$, its preimage $f^{-1}(U)\subseteq A(\!(t^{\mathbb{R}})\!)^{\oplus n}$ is open and hence contains some $t^{d}A[[t^{\mathbb{R}}]]^{\oplus n}$. It follows that $t^{d}M^{\circ}$ is contained in $U$.
Note that there is an ideal $(t^{>0})\subseteq A[[t^{\mathbb{R}}]]$ generated by the elements $t^{\epsilon}$ for all real numbers $\epsilon>0$. The quotient ring $A[[t^{\mathbb{R}}]]/(t^{>0})$ is naturally identified with $A$. Consider the $A$-submodule
$M_{0}=\{m^{\circ}\in M^{\circ}:F_{M}(m^{\circ})=m^{\circ}\},$
of $M^{\circ}$. We now show that the composition
$M_{0}\hookrightarrow M^{\circ}\twoheadrightarrow M^{\circ}/(t^{>0})M^{\circ}$
is an isomorphism of $A$-modules. For $m^{\circ},n^{\circ}\in M^{\circ}$, if $m^{\circ}-n^{\circ}\in t^{d}M^{\circ}$ then $F_{M}(m^{\circ})-F_{M}(n^{\circ})\in t^{\lambda d}M^{\circ}$, because $M^{\circ}$ is stable under $F_{M}$. This implies that for every $m^{\circ}\in M^{\circ}$, the limit
$\tilde{m}^{\circ}=\lim_{k\to\infty}F_{M}^{k}(m^{\circ})$
exists, because if $F_{M}(m^{\circ})-m^{\circ}\in t^{\epsilon}M^{\circ}$ then $F_{M}^{k+1}(m^{\circ})-F_{M}^{k}(m^{\circ})\in t^{\lambda^{k}\epsilon}M^{\circ}$. Since $\tilde{m}^{\circ}\in M_{0}$ and $\tilde{m}^{\circ}-m^{\circ}\in(t^{>0})M^{\circ}$, we conclude that the composition $M_{0}\to M^{\circ}/(t^{>0})M^{\circ}$ is surjective. Similarly, if $n_{0},m_{0}\in M_{0}$ satisfy $n_{0}-m_{0}\in(t^{>0})M^{\circ}$, then by iteratively applying $F_{M}$ we see that $n_{0}-m_{0}\in t^{d}M^{\circ}$ for all $d\in\mathbb{R}$, and hence $n_{0}=m_{0}$.
The isomorphism $M_{0}\cong M^{\circ}/(t^{>0})M^{\circ}$ in particular implies that $M_{0}$ is finitely generated as an $A$-module. Choose a finite set of generators $m_{1},\dotsc,m_{n}\in M_{0}$. Because any element in $1+(t^{>0})\subseteq A[[t^{\mathbb{R}}]]$ is invertible, Nakayama’s lemma implies that $m_{1},\dotsc,m_{n}$ generate $M^{\circ}$ as an $A[[t^{\mathbb{R}}]]$-module. It follows that
$A(\!(t^{\mathbb{R}})\!)^{\oplus n}\xrightarrow{(m_{1},\dotsc,m_{n})}M$
is surjective. We can further promote it to a surjective map of Novikov isocrystals
$(A(\!(t^{\mathbb{R}})\!)^{\oplus n},F^{\oplus n})\xrightarrow{(m_{1},\dotsc,m_{n})}(M,F_{M}).$
We now consider the kernel
$K=\ker(A(\!(t^{\mathbb{R}})\!)^{\oplus n}\to M).$
This is a closed $A(\!(t^{\mathbb{R}})\!)$-submodule of $(A^{\oplus n})(\!(t^{\mathbb{R}})\!)$ that is stable under $F^{\oplus n}$. Lemma 3.6 implies that there exists a $A$-submodule $K_{0}\subseteq A^{\oplus n}$ for which $K=K_{0}(\!(t^{\mathbb{R}})\!)$. On the other hand, $K$ has the subspace topology from $A(\!(t^{\mathbb{R}})\!)^{\oplus n}$, and therefore the $t$-adic topology on $K_{0}(\!(t^{\mathbb{R}})\!)$ agrees with the canonical topology. Hence Lemma 2.10 shows that $K_{0}$ is a finite projective $A$-module. At this point, we have
$(M,F_{M})$ $\cong(A(\!(t^{\mathbb{R}})\!)^{\oplus n},F^{\oplus n})/(K_{0}(\!(t^{\mathbb{R}})\!)^{\oplus n},\operatorname{id}_{K_{0}}\otimes F)$
$=((A^{\oplus n}/K_{0})(\!(t^{\mathbb{R}})\!),\operatorname{id}_{A^{\oplus n}/K_{0}}\otimes F).$
Finally, as in the proof of Lemma 2.10, $A^{\oplus n}/K_{0}$ being finitely presented over $A$ with $(A^{\oplus n}/K_{0})(\!(t^{\mathbb{R}})\!)$ being finite projective over $A(\!(t^{\mathbb{R}})\!)$ implies that $A^{\oplus n}/K_{0}$ is finite projective over $A$, by [Sta24, Theorem 08XD]. ∎
4. Proof of the main theorem
4.1. Isocrystal from a descent datum
In this subsection, we construct a functor $\mathsf{NovDD}(A,\mathbb{R})\to\mathsf{NovIsoc}(A)$. First note that for every finite projective module $M$ over $A(\!(t^{\mathbb{R}})\!)$, the diagram
[…]
is an equalizer. Indeed, this can be checked for $M=A(\!(t^{\mathbb{R}})\!)$ by flatness of $M$, in which case it is clear.
We define the $A$-algebra automorphisms
$F_{2}$ $\colon A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!);$ $t$ $\mapsto t,u\mapsto u^{\lambda},$
$F_{3}$ $\colon A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!);$ $t$ $\mapsto t,u\mapsto u,v\mapsto v^{\lambda}.$
For each $(M,\varphi)\in\mathsf{NovDD}(A)$, we define a $F_{2}$-semilinear bijection
[…]
Similarly, we can define a $F_{3}$-semilinear bijection
$F_{M,3}\colon\rho_{3}^{\ast}M\xrightarrow{\pi_{13}^{\ast}\varphi^{-1}}\rho_{1}^{\ast}M\xrightarrow{\operatorname{id}_{M}\otimes F_{3}}\rho_{1}^{\ast}M\xrightarrow{\pi_{13}^{\ast}\varphi}\rho_{3}^{\ast}M,$
which agrees with the other possible definition
$F_{M,3}\colon\rho_{3}^{\ast}M\xrightarrow{\pi_{23}^{\ast}\varphi^{-1}}\rho_{2}^{\ast}M\xrightarrow{\operatorname{id}_{M}\otimes F_{3}}\rho_{2}^{\ast}M\xrightarrow{\pi_{23}^{\ast}\varphi}\rho_{3}^{\ast}M,$
thanks to the cocycle condition. They fit in a commutative diagram
[…]
Both rows are equalizers, and all horizontal maps are $A(\!(t^{\mathbb{R}})\!)$-linear while the vertical maps are $F$-semilinear, if we use the ring homomorphisms $\pi_{2}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ and $\rho_{3}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$ to give every module a $A(\!(t^{\mathbb{R}})\!)$-module structure. Therefore we obtain an $F$-semilinear bijection $F_{M}\colon M\to M$, and this defines the functor
$\mathsf{NovDD}(A,\mathbb{R})\to\mathsf{NovIsoc}(A).$
Lemma 4.1.
Let $A$ be any ring, and assume that the descent datum $(M,\varphi)\in\mathsf{NovDD}(A,\mathbb{R})$ induces $F_{M}\colon M\to M$. Then the diagram
[…]
commutes, where both rows are equalizers.
Proof.
The commutation of the left square is the cocycle condition, and the commutation of the two right squares is clear. We now verify that the first row is an equalizer. Note that because $F_{M,2}\colon\pi_{2}^{\ast}M\to\pi_{2}^{\ast}M$ is $F_{2}$-semilinear and its restriction along $\pi_{2}^{\ast}\colon M\hookrightarrow\pi_{2}^{\ast}M$ is $F_{M}$, we have $F_{M,2}=F_{M}\otimes F_{2}$. Using the definition of $F_{M,2}$, we see that the first row being an equalizer is equivalent to
[…]
being an equalizer. Since all maps are $A(\!(t^{\mathbb{R}})\!)$-linear if we use $\pi_{1}^{\ast}$ to give $\pi_{1}^{\ast}M$ the structure of an $A(\!(t^{\mathbb{R}})\!)$-module, we may use finite projectivity of $M$ to reduce to
[…]
being an equalizer, which is now clear. The second row can similarly be seen to be an equalizer. ∎
Proposition 4.2.
For every ring $A$, the functor $\mathsf{NovDD}(A,\mathbb{R})\to\mathsf{NovIsoc}(A)$ is fully faithful.
Proof.
It suffices to show that for two descent data $(M,\varphi_{M}),(N,\varphi_{N})\in\mathsf{NovDD}(A,\mathbb{R})$ and an $A(\!(t^{\mathbb{R}})\!)$-linear map $f\colon M\to N$, if $f\circ F_{M}=F_{N}\circ f$ then $\pi_{2}^{\ast}f\circ\varphi_{M}=\varphi_{N}\circ\pi_{1}^{\ast}f$.
We consider the two diagrams from Lemma 4.1 corresponding to $(M,\varphi_{M})$ and $(N,\varphi_{N})$. Since $f\circ F_{M}=F_{N}\circ f$ and the right squares on the diagram only involve $F_{M},F_{N}$, the universal property of equalizers produces unique maps $g\colon M\to N$ and $h\colon\pi_{2}^{\ast}M\to\pi_{2}^{\ast}N$ fitting in the commutative diagram
[…]
We note that $g$ must be $A(\!(t^{\mathbb{R}})\!)$-linear, since all maps in the top face are $A(\!(t^{\mathbb{R}})\!)$-linear when we use $\pi_{1}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ for the $A(\!(t^{\mathbb{R}})\!)$-module structure. Next, from the commutativity of the top face we see that the bottom face commutes when we replace $h$ with $\pi_{2}^{\ast}g$. By uniqueness of the morphisms, this implies that $h=\pi_{2}^{\ast}g$. On the other hand, by comparing the left face with the top face we obtain $h=\pi_{2}^{\ast}f$, as both $h$ and $\pi_{2}^{\ast}f$ are $A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$-linear while the image of $\varphi_{M}\circ\pi_{1}^{\ast}$ generates $\pi_{2}^{\ast}M$. Since $M,N$ are both finite projective and $\pi_{2}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, is injective, it follows from $\pi_{2}^{\ast}f=\pi_{2}^{\ast}g$ that $f=g$. Unraveling the top face now gives $\pi_{2}^{\ast}f\circ\varphi_{M}=\varphi_{N}\circ\pi_{1}^{\ast}f$. ∎
4.2. The case of a reduced ring
We now prove Theorem 4.9 when $A$ is reduced and $\Gamma=\mathbb{R}$. The first step is to prove it when $A$ is an algebraically closed field.
Proposition 4.3.
Let $k$ be an algebraically closed field. Then the functor $\mathsf{Vect}(k)\to\mathsf{NovDD}(k,\mathbb{R})$ is an equivalence of categories.
Proof.
In view of Proposition 2.6, it suffices to show that the functor is essentially surjective. Then by Proposition 4.2, it is enough to check that the image of $\mathsf{NovDD}(k,\mathbb{R})\to\mathsf{NovIsoc}(k)$ is contained in the essential image of $\mathsf{Vect}(k)\to\mathsf{NovIsoc}(k)$.
Assume that $(M,\varphi)\in\mathsf{NovDD}(k,\mathbb{R})$ induces a Novikov isocrystal $(M,F_{M})\in\mathsf{NovIsoc}(k)$. By Proposition 3.5, there exists a filtration
$0=M_{0}\hookrightarrow M_{1}\hookrightarrow\dotsb\hookrightarrow M_{n}=M$
with the proper that each associated graded $M_{i}/M_{i-1}$ is isomorphic to $(k(\!(t^{\mathbb{R}})\!),c_{i}F)$ for some $c_{i}\in k^{\times}$.
We inductively prove that each $(M_{i},F_{M}|_{M_{i}})\in\mathsf{NovIsoc}(k)$ is in the essential image of $\mathsf{Vect}(k)\to\mathsf{NovIsoc}(k)$, where the base case $i=0$ is clear. Assume we have an isomorphism
$(M_{i-1},F_{M}|_{M_{i-1}})\cong(k(\!(t^{\mathbb{R}})\!),F)^{\oplus(i-1)}.$
Fixing a splitting of the short exact sequence $0\to M_{i-1}\to M_{i}\to M_{i}/M_{i-1}\to 0$, just as $k(\!(t^{\mathbb{R}})\!)$-vector spaces, we find an isomorphism $M_{i}\cong k(\!(t^{\mathbb{R}})\!)^{\oplus i}$ under which $F_{M}|_{M_{i}}$ corresponds to
$F_{M}|_{M_{i}}=\begin{pmatrix}I_{i-1}&C\\ 0&c_{i}\end{pmatrix}F,$
where $I_{i-1}$ denotes the identity matrix and $C=(c_{1},\dotsc,c_{i-1})^{T}\in k(\!(t^{\mathbb{R}})\!)^{\oplus(i-1)}$ is a column vector.
On the other hand, by Lemma 4.1, the map
$\operatorname{id}-F_{M}|_{M_{i}}\otimes F_{2}\colon\pi_{2}^{\ast}M_{i}\to\pi_{2}^{\ast}M_{i}$
has an $i$-dimensional kernel. Since the kernel when restricted to $\pi_{2}^{\ast}M_{i-1}$ is $(i-1)$-dimensional, there exist $x_{1},\dotsc,x_{i}\in k(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ with $x_{i}\neq 0$ satisfying
$x_{1}=F_{2}(x_{1})+(\pi_{2}^{\ast}c_{1})F_{2}(x_{i}),\;\dotsc,\;x_{i-1}=F_{2}(x_{i-1})+(\pi_{2}^{\ast}c_{i-1})F_{2}(x_{i}),\;x_{i}=c_{i}F_{2}(x_{i}).$
Because $c_{i}\in k^{\times}$ and $x_{i}\neq 0$, it follows from the last equation that $x_{i}\in k(\!(t^{\mathbb{R}})\!)^{\times}\subseteq k(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ and $c_{i}=1$. Then setting $y_{j}=x_{j}/x_{i}$, we obtain
$y_{1}=F_{2}(y_{1})+\pi_{2}^{\ast}c_{1},\quad\dotsb,\quad y_{i-1}=F_{2}(y_{i-1})+\pi_{2}^{\ast}c_{i-1}.$
For each $y_{j}=\sum_{\alpha\in I_{j}}a_{\alpha,j}t^{d_{\alpha,j}}u^{e_{\alpha,j}}$, we define
$z_{j}=\sum_{\alpha\in I_{j},d_{\alpha,j}=0}a_{\alpha,j}t^{e_{\alpha,j}}\in k(\!(t^{\mathbb{R}})\!).$
Then we have
$z_{1}=F(z_{1})+c_{1},\quad\dotsb,\quad z_{i-1}=F(z_{i-1})+c_{i-1}.$
Finally, we can write
$F_{M}|_{M_{i}}=\begin{pmatrix}I_{i-1}&C\\ 0&1\end{pmatrix}F=\begin{pmatrix}I_{i-1}&Z\\ 0&1\end{pmatrix}F\begin{pmatrix}I_{i-1}&Z\\ 0&1\end{pmatrix}^{-1},$
where $Z=(z_{1},\dotsc,z_{i-1})^{T}\in k(\!(t^{\mathbb{R}})\!)^{\oplus(i-1)}$. This shows that under this change of basis, we have $(M_{i},F_{M}|_{M_{i}})\cong(k(\!(t^{\mathbb{R}})\!),F)^{\oplus i}$. ∎
Proposition 4.4.
Let $A=\prod_{i\in I}k_{i}$ be a product of algebraically closed fields. Then the functor $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\mathbb{R})$ is an equivalence.
Proof.
By Proposition 2.6, it remains to check that the functor is essentially surjective. Let $(M,\varphi)\in\mathsf{NovDD}(A,\mathbb{R})$ be a descent datum. By Proposition 4.3, for each $i$ there exists a $k_{i}$-vector space $M_{0,i}$ and an isomorphism
$f_{i}\colon M\otimes_{A(\!(t^{\mathbb{R}})\!)}k_{i}(\!(t^{\mathbb{R}})\!)\cong M_{0,i}\otimes_{k_{i}}k_{i}(\!(t^{\mathbb{R}})\!),$
compatible with the descent data. Say $M$ is generated by $N$ elements. Then $M\otimes_{A(\!(t^{\mathbb{R}})\!)}k_{i}(\!(t^{\mathbb{R}})\!)$ has dimension at most $N$, and hence $\dim_{k_{i}}M_{0,i}\leq N$. This uniform bound implies that $M_{0}=\prod_{i\in I}M_{0,i}$ is finite projective over $A$. It remains to show that $(M,\varphi)\cong(M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!),\operatorname{id})$.
Taking the product of $f_{i}$, we obtain an isomorphism
$f=\prod_{i\in I}f_{i}\colon M\otimes_{A(\!(t^{\mathbb{R}})\!)}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)\cong M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!).$
Compatibility of each $f_{i}$ with descent data means that
$\pi_{2}^{\ast}f_{i}\circ(\varphi\otimes_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!))=\pi_{1}^{\ast}f_{i},$
and taking the product over all $i$ gives a commutative diagram
[…]
In particular, if we consider $M$ as an $A(\!(t^{\mathbb{R}})\!)$-submodule of $M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ via $f$, the images of
$\pi_{1}^{\ast}M$ $=M\otimes_{A(\!(t^{\mathbb{R}})\!),\pi_{1}^{\ast}}A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)\hookrightarrow M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!),$
$\pi_{2}^{\ast}M$ $=M\otimes_{A(\!(t^{\mathbb{R}})\!),\pi_{2}^{\ast}}A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)\hookrightarrow M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!),$
agree.
For every element $\alpha=(\alpha_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, define the support of $\alpha$ as $\operatorname{supp}(\alpha)=\bigcup_{i\in I}\operatorname{supp}(\alpha_{i})\subseteq\mathbb{R}$, where the support of $\alpha_{i}\in M_{0,i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ is defined as before. For $\beta=(\beta_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ we similarly define $\operatorname{supp}(\beta)=\bigcup_{i}\operatorname{supp}(\beta_{i})\subseteq\mathbb{R}$.
We now claim that for each $m\in M\subseteq M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$, the support $\operatorname{supp}(m)\subseteq\mathbb{R}$ is bounded below. Choose $m_{1},\dotsc,m_{n}\in M\subseteq M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ a finite collection of generators as an $A(\!(t^{\mathbb{R}})\!)$-module. We can find elements $a_{jk}\in A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ for $1\leq j,k\leq n$ satisfying
($\ast$) $\pi_{1}^{\ast}m_{j}=\sum_{k=1}^{n}a_{jk}\pi_{2}^{\ast}m_{k}\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!).$
Fix a sufficiently small constant $C\in\mathbb{R}$ satisfying the property that
$\bigcup_{j,k=1}^{n}\operatorname{supp}(a_{jk})\subseteq\{(x,y)\in\mathbb{R}^{2}:2x+y\geq C\}.$
For each $i\in I$ and $1\leq j\leq n$, the set $\operatorname{supp}(m_{j,i})$ has a minimal element, and hence $D_{i}=\min\bigcup_{j=1}^{n}\operatorname{supp}(m_{j,i})$ exists. From ($\ast$) we see that
$\operatorname{supp}(m_{j,i})\times\{0\}\subseteq\bigcup_{k=1}^{n}(\operatorname{supp}(a_{jk,i})+(\{0\}\times\operatorname{supp}(m_{k,i})))$
inside $\mathbb{R}^{2}$, where by $+$ we mean the Minkowski sum. By our assumptions, the right hand side is always contained in $\{(x,y):2x+y\geq C+D_{i}\}$. On the other hand, for some $1\leq j\leq n$, the left hand side contains $(D_{i},0)$. This shows that $2D_{i}\geq C+D_{i}$, and hence $D_{i}\geq C$. This shows that $\operatorname{supp}(m_{j,i})\subseteq[C,\infty)$ for all $i,j$, and taking the union over all $i$, we obtain $\operatorname{supp}(m_{j})\subseteq[C,\infty)$ for all $1\leq j\leq n$. For an arbitrary $m\in M$, we can now write $m=\sum_{j=1}^{n}a_{j}m_{j}$ for $a_{j}\in A(\!(t^{\mathbb{R}})\!)$, and since each $\operatorname{supp}(a_{j})$ is bounded below, we conclude that $\operatorname{supp}(m)$ is bounded below as well.
Recall that the images of $\pi_{1}^{\ast}M\to M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ and $\pi_{2}^{\ast}M\to M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ agree. For each element $\xi$ in this common image, we can write $\xi=\sum_{j=1}^{n}a_{j}\pi_{1}^{\ast}m_{j}$, where $a_{j}\in A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ and $m_{j}\in M$. Then
$\operatorname{supp}(\xi)\subseteq\bigcup_{j}(\operatorname{supp}(a_{j})+(\operatorname{supp}(m_{j})\times\{0\}))$
inside $\mathbb{R}^{2}$. Because each $\operatorname{supp}(m_{j})$ is bounded below, for every $C\in\mathbb{R}$, the intersection
$\operatorname{supp}(\xi)\cap\{(x,y)\in\mathbb{R}^{2}:x+y\leq C\}$
has only finitely many $y$-coordinate values. For an arbitrary $m\in M$, applying this to $\xi=\pi_{2}^{\ast}m$ shows that the intersection $\operatorname{supp}(m)\cap(-\infty,C]$ is finite. That is, we have
$M\subseteq M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!).$
It remains to show that this inclusion is an equality. We apply the entire argument to the dual descent datum $M^{\vee}$ and the isomorphisms
$(f_{i}^{-1})^{\vee}\colon M^{\vee}\otimes_{A(\!(t^{\mathbb{R}})\!)}k_{i}(\!(t^{\mathbb{R}})\!)\cong M_{0,i}^{\vee}\otimes_{k_{i}}k_{i}(\!(t^{\mathbb{R}})\!).$
Then we similarly obtain that the image of the composition
$M^{\vee}\hookrightarrow M^{\vee}\otimes_{A(\!(t^{\mathbb{R}})\!)}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)\xrightarrow{(f^{-1})^{\vee}}M_{0}^{\vee}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$
lies in $M_{0}^{\vee}\otimes_{A}A(\!(t^{\mathbb{R}})\!)$. Dualizing the resulting map induces an $A(\!(t^{\mathbb{R}})\!)$-linear homomorphism $M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)\to M$ whose base change to $\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ recovers $f^{-1}$. Hence the composition
$M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)\to M\hookrightarrow M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)$
is the identity map upon base changing to $\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$. On the other hand, both sides are finite projective and $A(\!(t^{\mathbb{R}})\!)\hookrightarrow\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ is injective, and thus the map is the identity before base change. Therefore $M=M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)$. ∎
Proposition 4.5.
Let $A$ be a reduced ring. Then the functor $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\mathbb{R})$ is an equivalence.
Proof.
In view of Proposition 2.6, it suffices to show that the functor is essentially surjective. Then by Proposition 4.2, it is enough to check that the image of $\mathsf{NovDD}(A,\mathbb{R})\to\mathsf{NovIsoc}(A)$ is contained in the essential image of $\mathsf{Vect}(A)\to\mathsf{NovIsoc}(A)$. Consider any $(M,F_{M})\in\mathsf{NovIsoc}(A)$ induced from $(M,\varphi_{M})\in\mathsf{NovDD}(A,\mathbb{R})$. Since $A$ is reduced, we can embed $A$ into a product of algebraically closed fields as
$A\hookrightarrow B=\prod_{\mathfrak{p}}\overline{\operatorname{Frac}(A/\mathfrak{p})},$
where $\mathfrak{p}$ ranges over all prime ideals of $A$. Base changing along this ring homomorphism induces the descent datum $(M,\varphi_{M})\otimes_{A(\!(t^{\mathbb{R}})\!)}B(\!(t^{\mathbb{R}})\!)$. By Proposition 4.4, this is contained in the essential image of $\mathsf{Vect}(B)\to\mathsf{NovDD}(B,\mathbb{R})$. Hence the associated isocrystal $(M,F_{M})\otimes_{A(\!(t^{\mathbb{R}})\!)}B(\!(t^{\mathbb{R}})\!)$ lies in the essential image of $\mathsf{Vect}(B)\to\mathsf{NovIsoc}(B)$. We can now apply Proposition 3.7 to conclude that $(M,F_{M})$ is in the essential image of $\mathsf{Vect}(A)\to\mathsf{NovIsoc}(A)$. ∎
4.3. The case of a general ring
We now prove Theorem 4.9 when $A$ is an arbitrary commutative ring and $\Gamma=\mathbb{R}$. We begin by studying the square-zero deformation theory of Novikov descent data.
Lemma 4.6.
Let $A$ be a commutative ring, and let $I\subseteq A$ be an ideal satisfying $I^{2}=0$.
  1. (1)
    Let $M,N$ be two finite projective $A$-modules together with an isomorphism $\bar{f}\colon M/IM\cong N/IN$ of $A/I$-modules. Then $\bar{f}$ lifts to an isomorphism $f\colon M\cong N$ of $A$-modules.
  2. (2)
    Let $\bar{M}$ be a finite projective $A/I$-module. Then there exists a finite projective $A$-module $M$ together with an isomorphism $M/IM\cong\bar{M}$.
Proof.
(1) Letting $P=\operatorname{Hom}_{A}(M,N)$, we observe that $P$ is again a finite projective $A$-module with $\bar{f}\in P/IP$. Lift $\bar{f}$ to any element $f\in P$. We now check that $f$ is indeed an isomorphism $f\colon M\to N$ because we can work locally so that both $M$ and $N$ are free, and then the determinant being a unit can be checked modulo $I$.
(2) Note that $\operatorname{Spec}A$ is homeomorphic to $\operatorname{Spec}A/I$. There exists a cover $\{U_{i}\}$ of $\operatorname{Spec}A\cong\operatorname{Spec}A/I$ consisting of affine opens over which $\bar{M}$ is free. We lift $\bar{M}|_{U_{i}}$ to $M_{i}$ on each $U_{i}$ separately, and use (1) to choose isomorphisms $M_{i}|_{U_{i}\cap U_{j}}\cong M_{j}|_{U_{i}\cap U_{j}}$ lifting the identity map on $\bar{M}|_{U_{i}\cap U_{j}}$. These isomorphisms need not satisfy the cocycle condition, but the induced $2$-cocycle takes values in
$\ker(\mathrm{GL}_{A}(M_{i}|_{U_{i}})\to\mathrm{GL}_{A/I}(\bar{M}|_{U_{i}}))\cong\operatorname{Hom}_{A/I}(\bar{M},I\otimes_{A/I}\bar{M})|_{U_{i}}.$
Since the Čech $H^{2}$ of quasi-coherent sheaves with respect to $\{U_{i}\}$ vanish on $\operatorname{Spec}A/I$ as each $U_{i}$ is affine, this is a coboundary, and hence we may modify the isomorphisms $M_{i}|_{U_{i}\cap U_{j}}\cong M_{j}|_{U_{i}\cap U_{j}}$ so that they do satisfy the cocycle condition. ∎
Lemma 4.7.
Let $A$ be a commutative ring, and let $I\subseteq A$ be an ideal satisfying $I^{2}=0$. Let $(M,\varphi)\in\mathsf{NovDD}(A,\mathbb{R})$ be a descent datum satisfying the property that its base change to $A/I$ lies in the essential image of $\mathsf{Vect}(A/I)\to\mathsf{NovDD}(A/I,\mathbb{R})$. Then $(M,\varphi)$ is in the essential image of $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\mathbb{R})$.
Proof.
By assumption, there exists a finite projective $A/I$-module $\bar{M}_{0}$ together with an isomorphism
$(M/I(\!(t^{\mathbb{R}})\!)M,\bar{\varphi})\cong(\bar{M}_{0}(\!(t^{\mathbb{R}})\!),\operatorname{id})$
in $\mathsf{NovDD}(A/I,\mathbb{R})$. Using Lemma 4.6, we choose a finite projective $A$-module $M_{0}$ with $M_{0}/IM_{0}\cong\bar{M}_{0}$. As $A(\!(t^{\mathbb{R}})\!)\twoheadrightarrow(A/I)(\!(t^{\mathbb{R}})\!)$ is also a square-zero thickening, we may use Lemma 4.6 again to choose an isomorphism
$M\cong M_{0}(\!(t^{\mathbb{R}})\!)$
lifting the isomorphism $M/I(\!(t^{\mathbb{R}})\!)M\cong\bar{M}_{0}(\!(t^{\mathbb{R}})\!)\cong M_{0}(\!(t^{\mathbb{R}})\!)/(IM_{0})(\!(t^{\mathbb{R}})\!)$. Under this isomorphism, we obtain a new cocycle
$\psi\colon M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)\xrightarrow{\cong}M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$
that reduces to the identity map on $(M_{0}/IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$.
It now suffices to show that $\psi$ is a coboundary, i.e., there exists an $A(\!(t^{\mathbb{R}})\!)$-linear automorphism $\xi\colon M_{0}(\!(t^{\mathbb{R}})\!)\cong M_{0}(\!(t^{\mathbb{R}})\!)$ for which $\psi=\pi_{2}^{\ast}\xi^{-1}\circ\pi_{1}^{\ast}\xi$. We first write
$\psi=\operatorname{id}+\psi_{+},\quad\psi\colon M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)\to(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!).$
Since $I$ squares to zero, the cocycle condition on $\psi$ translates to an additive cocycle condition
$\pi_{13}^{\ast}\psi_{+}=\pi_{23}^{\ast}\psi_{+}+\pi_{12}^{\ast}\psi_{+},$
as $A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$-linear maps $M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)\to(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$. Here, we note that $(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)=I(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$, and this justifies that $\pi_{23}^{\ast}\psi_{+}\circ\pi_{12}^{\ast}\psi_{+}=0$.
As $M_{0}$ is finite projective over $A$, there are canonical isomorphisms
$\operatorname{Hom}_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)}(M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!),(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!))$ $\cong\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!),$
$\operatorname{Hom}_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)}(M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!),(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!))$ $\cong\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!).$
Using this isomorphism, we identify $\psi_{+}$ as an element of $\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, and similarly the additive cocycle condition as an equation valued in the group $\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$. By considering the coefficients in the cocycle condition, we readily see that $\psi_{+}$ is supported on ({0})×{0}{0}×($(\mathbb{R}\setminus\{0\})\times\{0\}\cup\{0\}\times(\mathbb{R}\setminus\{0\})$, and moreover the coefficient of $t^{d}$ agrees with the coefficient of $u^{d}$ for each $d\in\mathbb{R}\setminus\{0\}$. This allows us to write
$\psi_{+}=\pi_{1}^{\ast}\xi_{+}-\pi_{2}^{\ast}\xi_{+},\quad\xi_{+}\in\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}})\!).$
Regarding $\xi_{+}$ as a $A(\!(t^{\mathbb{R}})\!)$-linear map $M_{0}(\!(t^{\mathbb{R}})\!)\to(IM_{0})(\!(t^{\mathbb{R}})\!)$, we now define $\xi=\operatorname{id}+\xi_{+}$. It follows that
$\psi=\operatorname{id}+\pi_{1}^{\ast}\xi_{+}-\pi_{2}^{\ast}\xi_{+}=(\operatorname{id}-\pi_{2}^{\ast}\xi_{+})\circ(\operatorname{id}+\pi_{1}^{\ast}\xi_{+})=\pi_{2}^{\ast}\xi^{-1}\circ\pi_{1}^{\ast}\xi.$
This finishes the proof. ∎
Proposition 4.8.
Let $A$ be a commutative ring. Then the functor $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\mathbb{R})$ is an equivalence.
Proof.
By Proposition 2.6, it suffices to show that the functor is essentially surjective. Consider a Novikov descent datum $(M,\varphi)\in\mathsf{NovDD}(A,\mathbb{R})$, and let $(M,F_{M})$ be the Novikov isocrystal attached to it. Let $I\subseteq A$ be the nilradical so that $B=A/I$ is reduced. By Proposition 4.5, there exists a finite projective $B$-module $N_{0}$ for which
$(M\otimes_{A(\!(t^{\mathbb{R}})\!)}B(\!(t^{\mathbb{R}})\!),F_{M}\otimes F)\cong(N_{0}\otimes_{B}B(\!(t^{\mathbb{R}})\!),\operatorname{id}_{N_{0}}\otimes F)$
as Novikov isocrystals over $B$.
Choose a finite set of generators $n_{1},\dotsc,n_{k}\in N_{0}$ as $B$-modules. Because $A(\!(t^{\mathbb{R}})\!)\to B(\!(t^{\mathbb{R}})\!)$ is surjective, there is also a surjection
$M\to M\otimes_{A(\!(t^{\mathbb{R}})\!)}B(\!(t^{\mathbb{R}})\!)\cong N_{0}(\!(t^{\mathbb{R}})\!).$
Lift the elements $n_{1},\dotsc,n_{k}\in N_{0}\subseteq N_{0}(\!(t^{\mathbb{R}})\!)$ to elements $m_{1},\dotsc,m_{k}\in M$. Since elements of $1+I(\!(t^{\mathbb{R}})\!)$ are invertible in $A(\!(t^{\mathbb{R}})\!)$, the elements $m_{1},\dotsc,m_{k}$ generate $M$ as a $A(\!(t^{\mathbb{R}})\!)$-module by Nakayama’s lemma. This implies that if we set
$M^{\circ}=A[[t^{\mathbb{R}}]]m_{1}+\dotsb+A[[t^{\mathbb{R}}]]m_{k}$
then $M^{\circ}[t^{-1}]=M$.
For each $1\leq i\leq k$, we have
$F_{M}(m_{i})-m_{i}\in\ker(M\to M\otimes_{A(\!(t^{\mathbb{R}})\!)}B(\!(t^{\mathbb{R}})\!))=M\otimes_{A(\!(t^{\mathbb{R}})\!)}I(\!(t^{\mathbb{R}})\!).$
That is, we have
$F_{M}(m_{i})\in m_{i}+I(\!(t^{\mathbb{R}})\!)m_{1}+I(\!(t^{\mathbb{R}})\!)m_{2}+\dotsb+I(\!(t^{\mathbb{R}})\!)m_{k}.$
Separating out the monomials with non-positive degree, we can further write
$F_{M}(m_{i})-m_{i}-\sum_{j\in I_{i}}a_{i,j}t^{d_{i,j}}m_{l_{i,j}}\in\sum_{i=1}^{k}(t^{>0})I[[t^{\mathbb{R}}]]m_{i}\subseteq(t^{>0})M^{\circ},$
where $I_{i}$ is finite and $a_{i,j}\in I$. That is, if we consider the finitely generated ideal $J\subseteq I$ generated by $a_{i,j}$ over all $1\leq i\leq k$ and $j\in I_{i}$, we obtain
$F_{M}(m_{i})-m_{i}\in(t^{>0})M^{\circ}$
modulo $J(\!(t^{\mathbb{R}})\!)$. Defining $C=A/J$, so that $A\twoheadrightarrow C\twoheadrightarrow B$, we conclude that the Novikov isocrystal
$(M\otimes_{A(\!(t^{\mathbb{R}})\!)}C(\!(t^{\mathbb{R}})\!),F_{M}\otimes F)$
has the property that the $C[[t^{\mathbb{R}}]]$-submodule of $M\otimes_{A(\!(t^{\mathbb{R}})\!)}C(\!(t^{\mathbb{R}})\!)$ generated by the images of $m_{1},\dotsc,m_{k}$ satisfies the hypotheses of Proposition 3.8. This shows that the base change of $(M,F_{M})$ to $C$ lies in the essential image of $\mathsf{Vect}(C)\to\mathsf{NovIsoc}(C)$, and by Proposition 4.2, it also follows that the base change of $(M,\varphi)$ to $C$ lies in the essential image of $\mathsf{Vect}(C)\to\mathsf{NovDD}(C,\mathbb{R})$.
On the other hand, because all elements of $J$ are nilpotent and $J$ is finitely generated, some finite power of $J$ is the zero ideal. This allows us to iteratively apply Lemma 4.7 to conclude that $(M,\varphi)$ lies in the essential image over $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\mathbb{R})$. ∎
4.4. Restricting the exponents
We finally remove the hypothesis that $\Gamma=\mathbb{R}$. Note there is a natural functor $\mathsf{NovDD}(A,\Gamma)\to{\mathsf{NovDD}(A,\mathbb{R})}$ given by tensoring along the ring homomorphism $A(\!(t^{\Gamma})\!)\to{A(\!(t^{\mathbb{R}})\!)}$.
Theorem 4.9.
Let $A$ be a commutative ring, and let $\Gamma\subseteq\mathbb{R}$ be an additive submonoid. Then the functor $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma)$ is an equivalence of categories.
Proof.
By Proposition 2.6, it is enough to prove essential surjectivity. Let $(M,\varphi)\in\mathsf{NovDD}(A,\Gamma)$ be a Novikov descent datum. By Proposition 4.8, there exists a finite projective $A$-module $M_{0}$ together with an isomorphism
$f\colon(M\otimes_{A(\!(t^{\Gamma})\!)}{A(\!(t^{\mathbb{R}})\!)},\varphi\otimes_{A(\!(t^{\Gamma},u^{\Gamma})\!)}{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)})\xrightarrow{\cong}({M_{0}(\!(t^{\mathbb{R}})\!)},\operatorname{id}_{M_{0}}\otimes{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)})$
in ${\mathsf{NovDD}(A,\mathbb{R})}$. This in particular allows us to regard $M$ as an $A(\!(t^{\Gamma})\!)$-submodule of ${M_{0}(\!(t^{\mathbb{R}})\!)}$ with the property that the images of
$M\otimes_{A(\!(t^{\Gamma})\!),\pi_{1}^{\ast}}A(\!(t^{\Gamma},u^{\Gamma})\!)\hookrightarrow{M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)},\quad M\otimes_{A(\!(t^{\Gamma})\!),\pi_{2}^{\ast}}A(\!(t^{\Gamma},u^{\Gamma})\!)\hookrightarrow{M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)}$
agree.
Consider any element $x$ in the common image. Since it can be written as a finite sum
$x=\sum_{i=1}^{n}(\pi_{1}^{\ast}m_{i})a_{i},\quad m_{i}\in M,\quad a_{i}\in A(\!(t^{\Gamma},u^{\Gamma})\!),$
we see that $\operatorname{supp}(x)$ must be contained in ×$\mathbb{R}\times\Gamma$. For every element $m\in M$, we in particular have $\operatorname{supp}(\pi_{2}^{\ast}m)\subseteq\mathbb{R}\times\Gamma$, and hence $\operatorname{supp}(m)\subseteq\Gamma$. This shows that
$M\subseteq M_{0}(\!(t^{\Gamma})\!),$
and it now suffices to prove that this inclusion is an equality.
We achieve this by running the same argument for the dual descent datum. Write $g\colon M\hookrightarrow M_{0}(\!(t^{\Gamma})\!)$ for the inclusion above. If we consider the linear dual of $f$, we obtain an isomorphism
$(f^{\vee})^{-1}\colon M^{\vee}\otimes_{A(\!(t^{\Gamma})\!)}{A(\!(t^{\mathbb{R}})\!)}\xrightarrow{\cong}{M_{0}^{\vee}(\!(t^{\mathbb{R}})\!)}$
compatible with descent data. By the argument above, there exists an inclusion $h\colon M^{\vee}\hookrightarrow M_{0}^{\vee}(\!(t^{\Gamma})\!)$ that recovers $(f^{\vee})^{-1}$ upon base changing along $A(\!(t^{\Gamma})\!)\to{A(\!(t^{\mathbb{R}})\!)}$. This shows that the composition
[…]
is the identity map upon base changing along $A(\!(t^{\Gamma})\!)\to{A(\!(t^{\mathbb{R}})\!)}$. Because both sides are finite projective and $A(\!(t^{\Gamma})\!)\hookrightarrow{A(\!(t^{\mathbb{R}})\!)}$ is injective, the composition $g\circ h^{\vee}$ is the identity map. In particular, $g$ is surjective, and therefore a bijection. ∎
5. Application to perfectoid geometry
We now record some consequences of our descent result. We will use Theorem 4.9 with the exponent submonoid $\Gamma=\mathbb{Z}[p^{-1}]$. It is customary to write
${A(\!(t^{\Gamma})\!)}=A(\!(t^{{1/p^{\infty}}})\!)$
in the setting of perfectoid geometry.
Let $\mathsf{Perf}$ be the category of affinoid perfectoid spaces in characteristic $p$, equipped with the v-topology. For a perfect $\mathbb{F}_{p}$-algebra $A$, we consider the v-sheaf
$\operatorname{Spd}A\colon\mathsf{Perf}^{\mathrm{op}}\to\mathsf{Set};\quad\operatorname{Spa}(R,R^{+})\mapsto\operatorname{Hom}(A,R^{+}).$
Similarly for $B$ a perfect (hence perfectoid) complete Tate $\mathbb{F}_{p}$-algebra, we define the v-sheaf
$\operatorname{Spd}B=\operatorname{Spd}(B,B^{\circ})\colon\mathsf{Perf}^{\mathrm{op}}\to\mathsf{Set};\quad\operatorname{Spa}(R,R^{+})\mapsto\operatorname{Hom}_{\mathrm{cts}}((B,B^{\circ}),(R,R^{+})).$
We refer the reader to [SW20] for the detailed definitions of the objects appearing in this section.
Definition 5.1.
For a v-sheaf $X$, we define a vector bundle on $X$ to be an association to each $(R,R^{+})$-point $\xi\in X(R,R^{+})$ a locally finite free sheaf $\mathscr{V}_{\xi}$ on $\operatorname{Spa}(R,R^{+})$, together with the data of, for each map $f\colon\operatorname{Spa}(S,S^{+})\to\operatorname{Spa}(R,R^{+})$ and $\xi\in X(R,R^{+})$, an isomorphism $f^{\ast}\mathscr{V}_{\xi}\cong\mathscr{V}_{f^{\ast}\xi}$, satisfying the condition that for every
[…]
and $\xi\in X(R,R^{+})$, the three identifications between $g^{\ast}f^{\ast}\mathscr{V}_{\xi}$, $g^{\ast}\mathscr{V}_{f^{\ast}\xi}$, and $\mathscr{V}_{g^{\ast}f^{\ast}\xi}$ commute.
Corollary 5.2.
Let $A$ be a perfect $\mathbb{F}_{p}$-algebra. Then the natural functor
{vector bundles on $\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
is an exact tensor equivalence.
Proof.
As the natural map
$X=\operatorname{Spd}(A(\!(t^{{1/p^{\infty}}})\!),A[[t^{{1/p^{\infty}}}]])\to\operatorname{Spd}A=Y$
is a v-cover, by [SW20, Lemma 17.1.8] the category of vector bundles on $\operatorname{Spd}A$ can be identified with the category of pairs $(\mathscr{V},\varphi)$, where $\mathscr{V}$ is a vector bundle on $X$ and $\varphi$ is an isomorphism of vector bundles over $X\times_{Y}X$ satisfying a cocycle condition over $X\times_{Y}X\times_{Y}X$. By [KL15, Theorem 2.7.7] and again [SW20, Lemma 17.1.8], the vector bundle $\mathscr{V}$ corresponds to a finite projective module $M$ over $A(\!(t^{{1/p^{\infty}}})\!)$.
Next, we identify
$X\times_{Y}X=\bigcup_{n=2}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr{\rangle}[1/tu].$
If we denote by $\pi_{1},\pi_{2}\colon X\times_{Y}X\to X$ the two projection maps, then $\varphi$ is an isomorphism of vector bundles
$\varphi\colon\pi_{1}^{\ast}\mathscr{V}\xrightarrow{\cong}\pi_{2}^{\ast}\mathscr{V}$
over $X\times_{Y}X$. Because
$H^{0}(X\times_{Y}X,\mathscr{O}_{X\times_{Y}X})=\bigcap_{n=2}^{\infty}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr{\rangle}[1/tu]=A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!),$
we see that $\varphi$ corresponds to an isomorphism
$\varphi\colon M\otimes_{A(\!(t^{{1/p^{\infty}}})\!),\pi_{1}^{\ast}}A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!)\xrightarrow{\cong}M\otimes_{A(\!(t^{{1/p^{\infty}}})\!),\pi_{2}^{\ast}}A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!).$
Finally the cocycle condition is defined on
$X\times_{Y}X\times_{Y}X=\bigcup_{n=3}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{uv},\frac{u^{n}}{tv},\frac{v^{n}}{tu}\Bigr{\rangle}[1/tuv],$
and hence be checked after tensoring to
$H^{0}(X\times_{Y}X\times_{Y}X,\mathscr{O}_{X\times_{Y}X\times_{Y}X})=A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}})\!).$
This shows that a vector bundle on $\operatorname{Spd}A$ corresponds to a Novikov descent datum for $A$. Applying Theorem 4.9 now gives the desired result. ∎
We now compare the finite étale sites of $\operatorname{Spec}A$ and $\operatorname{Spd}A$.
Definition 5.3 ([Sch22, Definition 10.1]).
We say that a map $f\colon X\to Y$ of v-sheaves is finite étale when for every map $\operatorname{Spa}(R,R^{+})\to Y$, the fiber product $\operatorname{Spa}(R,R^{+})\times_{Y}X$ is representable by an affinoid perfectoid $\operatorname{Spa}(S,S^{+})$ where $R\to S$ is finite étale and $S^{+}$ is the integral closure of the image of $R^{+}$.
Corollary 5.4.
Let $A$ be a perfect $\mathbb{F}_{p}$-algebra. Then the natural functor
[…]
is an equivalence.
Proof.
We note that a finite étale $R$-algebra is necessarily a finite projective $R$-module. By Corollary 5.2, we have an equivalence between the category of vector bundles on $\operatorname{Spec}A$ with an algebra structure and the category of vector bundles on $\operatorname{Spd}A$ with an algebra structure. It remains to check that the condition of being étale agrees on both sides.
If the finite $A$-algebra $B$ is étale, it is clear that the corresponding $R$-algebra $R\otimes_{A}B$ is also étale over $R$ for every $A\to R^{+}$. In the other direction we note that if $A(\!(t^{{1/p^{\infty}}})\!)\to B\otimes_{A}A(\!(t^{{1/p^{\infty}}})\!)$ is finite étale, then so is $A\to B$ by [Sta24, Theorem 08XE], as the inclusion $A\to A(\!(t^{{1/p^{\infty}}})\!)$ splits as a map of $A$-modules. ∎

### v2

@@
 Let $A$ be a commutative ring and let $\Gamma\subseteq\mathbb{R}$ be an additive submonoid. Then the category of descent data of finite projective modules along the semi-cosimplicial diagram
-$A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\rightrightarrows\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!)$
+$A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\raisebox{1.70717pt}{$\rightrightarrows$}\hskip-10.00002pt\raisebox{-1.70717pt}{$\rightrightarrows$}A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!)$
 is equivalent to the category of finite projective $A$-modules.
 Remark 1.2.
-In general, the ring $A(\!(t^{\Gamma},u^{\Gamma})\!)$ is much larger than the tensor product $A(\!(t^{\Gamma})\!)\otimes_{A}A(\!(t^{\Gamma})\!)$, which is usually where the isomorphism of the descent datum is defined.
+In general, the ring $A(\!(t^{\Gamma},u^{\Gamma})\!)$ is much larger than the tensor product $A(\!(t^{\Gamma})\!)\otimes_{A}A(\!(t^{\Gamma})\!)$, which is usually where the isomorphism of the descent datum is defined over.
 As a consequence, standard descent results for finite projective modules are inapplicable to this setting.
+Remark 1.3.
+This may heuristically be interpreted as an analytic-to-discrete descent result for adic spaces.
+For example, let $k$ be a field, consider the field k((t))k(\!(t)\!) of Laurent series with the $t$-adic topology, and take its adic spectrum $X=\operatorname{Spa}(k(\!(t)\!),k[[t]])$.
+Then we have
+$Y=X\times_{\operatorname{Spa}(k,k)}X=\bigcup_{n\geq 1}\operatorname{Spa}\Bigl(k\Bigl[t,u,\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr]^{\wedge}\Bigl[\frac{1}{tu}\Bigr],k\Bigl[t,u,\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr]^{\wedge}\Bigr)$
+in the category of adic spaces, where we use the $(tu)$-adic topology for the completion.
+The adic space $Y$ can also be regarded as the punctured open unit disc over k((t))k(\!(t)\!), and we have $H^{0}(Y,\mathscr{O}_{Y})=k(\!(t,u)\!)$.
+Then Theorem 1.1 implies that the category of vector bundles on $X$ equipped with descent data over $Y$ is equivalent to the category of vector bundles on $\operatorname{Spa}(k,k)$.
 The theorem is motivated by question in perfectoid geometry, namely that of descending a vector bundle from $\operatorname{Spd}A$ to a vector bundle on $\operatorname{Spec}A$, where $A$ is a discrete perfect ring in characteristic $p$.
 Using Theorem 1.1 with $\Gamma=\mathbb{Z}[p^{-1}]$, we deduce the following result as an immediate consequence.
-Theorem 1.3 (Corollary 5.2, 5.4).
+We refer the reader to Section 5 for the definitions of the objects involved.
+Theorem 1.4 (Corollary 5.3, 5.5).
 Let $A$ be a perfect $\mathbb{F}_{p}$-algebra. Then the natural functor
-{vector bundles on $\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
+$\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
 is an exact tensor equivalence. Similarly, the natural functor
@@
 It is not true in general that a v-sheaf étale over $\operatorname{Spd}A$ descends to a scheme étale over $\operatorname{Spec}A$.
-For example, consider the perfect $\mathbb{F}_{p}$-algebra $A=\mathbb{F}_{p}[t^{1/p^{\infty}}]$.
-The associated v-sheaf is $(\operatorname{Spd}A)(R,R^{+})=R^{+}$, and admits a v-subsheaf $X(R,R^{+})=R^{\circ\circ}$.
+Consider the following example from [SW20, Example 18.2.1].
+Take the perfect $\mathbb{F}_{p}$-algebra $A=\mathbb{F}_{p}[t^{1/p^{\infty}}]$, which has associated v-sheaf $(\operatorname{Spd}A)(R,R^{+})=R^{+}$ that admits a v-subsheaf $X(R,R^{+})=R^{\circ\circ}$.
 Then $X\to\operatorname{Spd}A$ is an open embedding, hence étale, but does not correspond to an étale $A$-scheme.
@@
 On the other hand, the second part of the theorem remains true by [Sch22, Lemma 15.6].
-The second part of the theorem, Corollary 5.4, will be used in a future work [DKvHZ24] to construct integral models of Hodge type Shimura varieties at parahoric level.
-Specifically, it allows one to construct a finite étale cover of the special fiber a Shimura variety from the finite étale cover of the associated v-sheaf.
+The second part of Theorem 1.4 is used in [DKvHZ24] to construct integral models of Hodge type Shimura varieties at parahoric level.
+Specifically, it allows one to construct a finite étale cover of the special fiber of a Shimura variety from the finite étale cover of the associated v-sheaf.
 Remark 1.7.
-Ian Gleason announced via private communication an alternative proof of Corollary 5.4, which works by comparing the category of étale $\mathbb{F}_{\ell}$-local systems on $\operatorname{Spd}A$ and $\operatorname{Spec}A$.
+Ian Gleason announced via private communication an alternative proof of Corollary 5.5, i.e., the finite étale part of Theorem 1.4, which works by comparing the category of étale $\mathbb{F}_{\ell}$-local systems on $\operatorname{Spd}A$ and $\operatorname{Spec}A$.
 1.1. Outline of proof
 We sketch the main ideas that go into the proof of Theorem 1.1.
-In Section 3, we introduce the notion of a Novikov isocrystal, which somewhat easier to control than a descent datum as it involves only one variable.
+In Section 3, we introduce the notion of a Novikov isocrystal, which is somewhat easier to control than a descent datum as it involves only one variable.
 In Subsection 4.1, we construct a functor from the category of Novikov descent data for $\Gamma=\mathbb{R}$ to the category of Novikov isocrystals, and prove that this functor is always fully faithful.
@@
 We first take care the case when $A$ is an algebraically closed field by using the structure theory of Novikov isocrystals developed in Section 3.
-Next, we check that when $A$ is an arbitrary product of algebraically closed fields, effectivity over $A$ follows from that over each field.
+Next, we check that when $A$ is an arbitrary product of algebraically closed fields, effectivity over $A$ from that over each field.
 This allows to deduce effectivity for a general reduced ring $A$ by embedding $A$ into the product of the algebraic closures of its residue fields.
@@
 I would also like to thank my advisor Richard Taylor for his encouragements and diligence during our meetings, where we carefully went through parts of the argument.
-I also would like to thank Brian Conrad, Ian Gleason, and Kiran Kedlaya for helpful comments and suggestions.
+I also would like to thank Brian Conrad, Ian Gleason, Kiran Kedlaya, and the anonymous referee for helpful comments and suggestions.
 2. Novikov rings
@@
 Remark 2.2.
+We may also describe the ring $A(\!(t^{\Gamma})\!)$ and the module $M(\!(t^{\Gamma})\!)$ as follows.
+We first consider the polynomial ring $A[t^{\Gamma}]$ and the module $M[t^{\Gamma}]$ where elements of $A[t^{\Gamma}]$ are finite sums of monomials $a_{i}t^{d_{i}}$.
+We endow $A[t^{\Gamma}]$ with the topology where $\bigoplus_{c<d\in\Gamma}At^{d}\subseteq A[t^{\Gamma}]$ for $c\in\mathbb{R}$ form a neighborhood basis of 0, and similarly for $M[t^{\Gamma}]$.
+Then $A(\!(t^{\Gamma})\!)$ and $M(\!(t^{\Gamma})\!)$ are the completions of the topological ring $A[t^{\Gamma}]$ and the topological module $M[t^{\Gamma}]$.
+Remark 2.3.
 Such rings naturally appear as coefficient rings in Lagrangian Floer theory [FOOO09], and are named after Novikov for his use of a formal variable in the coefficient ring of a homology theory to keep track of topological invariants [Nov81].
-We note that the functor $M\mapsto M(\!(t^{\Gamma})\!)$ sending $A$-modules to $A(\!(t^{\Gamma})\!)$-modules commutes with limits and finite colimits.
-In particular, when $M$ is a finitely presented $A$-module, the natural map
-$M\otimes_{A}A(\!(t^{\Gamma})\!)\to M(\!(t^{\Gamma})\!)$
-is an isomorphism. As a consequence, for any other $A$-module $N$ there is a canonical isomorphism
-$\operatorname{Hom}_{A(\!(t^{\Gamma})\!)}(M(\!(t^{\Gamma})\!),N(\!(t^{\Gamma})\!))\cong\operatorname{Hom}_{A}(M,N)(\!(t^{\Gamma})\!)$
-of $A(\!(t^{\Gamma})\!)$-modules.
 To formulate our descent problem, we define a version of the Novikov ring with multiple variables.
@@
 where we note that the definitions of $A(\!(t^{\Gamma})\!)$ and $M(\!(t^{\Gamma})\!)$ agree with that of Definition 2.1.
+Example 2.5.
+When $\Gamma=\mathbb{Z}_{\leq 0}$, we have $A(\!(t^{\Gamma},u^{\Gamma})\!)=A[t^{-1},u^{-1}]$, and when $\Gamma=\mathbb{Z}_{\geq 0}$, we have $A(\!(t^{\Gamma},u^{\Gamma})\!)=A[[t,u]]$.
+When $\Gamma=\mathbb{Z}$, the ring A((t,u))A(\!(t,u)\!) is in general much larger than the ring $A[[t,u]]{}[(tu)^{-1}]$ since it contains convergent power series such as
+$\sum_{n=0}^{\infty}t^{-n}u^{n^{2}}\in A(\!(t,u)\!).$
+On the other hand, even if $A$ is a field, the ring A((t,u))A(\!(t,u)\!) is not a field because tut-u is not invertible.
+Remark 2.6.
+As in Remark 2.2, we may construct the Novikov rings $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ as a completion of the polynomial ring $A[t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma}]$.
+Here, we given $A[t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma}]$ the topology generated by additive cosets of
+$\bigoplus_{d_{1},\dotsc,d_{n}\in\Gamma,\sum_{j}s_{j}d_{j}\geq C}At_{1}^{d_{1}}\dotsm t_{n}^{d_{n}}$
+where $s_{1},\dotsc,s_{n}\in\mathbb{R}_{>0}$ and $C\in\mathbb{R}$.
+This is indeed a topological ring, and its completion is $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$.
+One can similarly define a topological $A[t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma}]$-module $M[t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma}]$ whose completion is $M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$.
+In particular, it follows from this discussion that $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ indeed has a natural ring structure, and $M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ indeed has a natural structure of a module over it.
+Remark 2.7.
+In the case when $\Gamma\subseteq\mathbb{R}$ is a nontrivial subgroup, there is a more geometric interpretation.
+Fix an element $0<\gamma\in\Gamma$, and consider the ring $A[[t_{1}^{\Gamma_{\geq 0}},\dotsc,t_{n}^{\Gamma_{\geq 0}}]]$, which is completion of $A[t_{1}^{\Gamma_{\geq 0}},\dotsc,t_{n}^{\Gamma_{\geq 0}}]$ with respect to the $(t_{1}^{\gamma},\dotsc,t_{n}^{\gamma})$-adic topology.
+This defines an affinoid pre-adic space
+$X=\operatorname{Spa}(A[[t_{1}^{\Gamma_{\geq 0}},\dotsc,t_{n}^{\Gamma_{\geq 0}}]],A[[t_{1}^{\Gamma_{\geq 0}},\dotsc,t_{n}^{\Gamma_{\geq 0}}]]).$
+The locus on which $\lvert t_{1}\rvert,\dotsc,\lvert t_{n}\rvert>0$ is the union of the rational localizations
+$\operatorname{Spa}(R_{m},R_{m}^{+})$ $=\{\lvert-\rvert\in X:\lvert t_{j}\rvert>\lvert t_{i}\rvert^{m}>0\text{ for all }1\leq i,j\leq n\},$
+[…]
+[…]
+Lemma 2.8.
+The functor from $A$-modules to $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$-modules sending $M\mapsto M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ is exact.
+Proof.
+Given surjective $A$-module homomorphism $f\colon M_{1}\twoheadrightarrow M_{2}$, the induced map
+$f(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)\colon M_{1}(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)\to M_{2}(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$
+surjective because we can lift each nonzero coefficient individually.
+The kernel $\ker f(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ consists of those series whose coefficients all lie in $\ker f$, and hence is $(\ker f)(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$.
+∎
+In particular, when $M$ is a finitely presented $A$-module, the natural map
+$M\otimes_{A}A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)\to M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$
+is an isomorphism. As a consequence, for any other $A$-module NN there is a canonical isomorphism
+$\operatorname{Hom}_{A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)}(M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!),N(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!))\cong\operatorname{Hom}_{A}(M,N)(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$
+of $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$-modules.
+2.1. Novikov descent data
 We define $A$-algebra homomorphisms
-$\pi_{1}^{\ast}$ $\colon A(\!(t^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma})\!);$ $t$ t,maps-$\mapsto t,$
+$\pi_{1}^{\ast}$ $\colon A(\!(t^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma})\!);$ $t$ $\mapsto t,$
 $\pi_{2}^{\ast}$ $\colon A(\!(t^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma})\!);$ $t$ $\mapsto u,$
@@
 They form an augmented semi-cosimplicial diagram of $A$-algebras
-$A\to A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\rightrightarrows\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!).$
+$A\to A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\raisebox{1.67007pt}{$\rightrightarrows$}\hskip-10.00002pt\raisebox{-1.67007pt}{$\rightrightarrows$}A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!).$
 We will also use the notation $\rho_{1}^{\ast}=\pi_{12}^{\ast}\circ\pi_{1}^{\ast}=\pi_{13}^{\ast}\circ\pi_{1}^{\ast}$ and similarly $\rho_{2}^{\ast}=\pi_{12}^{\ast}\circ\pi_{2}^{\ast}=\pi_{23}^{\ast}\circ\pi_{1}^{\ast}$ and $\rho_{3}^{\ast}=\pi_{13}^{\ast}\circ\pi_{2}^{\ast}=\pi_{23}^{\ast}\circ\pi_{2}^{\ast}$.
-For a ring homomorphism $f^{\ast}\colon A\to B$ and an $A$-module $M$, we will also denote $f^{\ast}M=M\otimes_{A,f^{\ast}}B$, and abuse notation to write $f^{\ast}\colon M\to f^{\ast}M$ for the map $m\mapsto m\otimes 1$.
+For a ring homomorphism $f^{\ast}\colon A\to B$ and an $A$-module $M$, we will also denote $f^{\ast}M=M\otimes_{A,f^{\ast}}B$, and write $f^{\ast}\colon M\to f^{\ast}M$ for the map $m\mapsto m\otimes 1$ by making an abuse of notation.
 Definition 2.9.
@@
 $\varphi\colon\pi_{1}^{\ast}M\xrightarrow{\cong}\pi_{2}^{\ast}M$
-is a $A(\!(t^{\Gamma},u^{\Gamma})\!)$-linear automorphism satisfying the cocycle condition
-[…]
-We denote by $\mathsf{NovDD}(A,\Gamma)$ the category of Novikov descent data.
+is an $A(\!(t^{\Gamma},u^{\Gamma})\!)$-linear automorphism satisfying the cocycle condition, which states that the diagram
+$\rho_{1}^{\ast}M$$\rho_{3}^{\ast}M$$\rho_{2}^{\ast}M.$$\scriptstyle{\pi_{13}^{\ast}\varphi}$$\scriptstyle{\pi_{12}^{\ast}\varphi}$$\scriptstyle{\pi_{23}^{\ast}\varphi}$
+commutes. We denote by $\mathsf{NovDD}(A,\Gamma)$ the category of Novikov descent data.
 Denote by $\mathsf{Vect}(A)$ the category of finite projective $A$-modules. There is a natural functor
-$\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma);\quad M_{0}\mapsto(M_{0}\otimes_{A}A(\!(t^{\Gamma})\!),\operatorname{id}_{M_{0}\otimes_{A}A(\!(t^{\Gamma},u^{\Gamma})\!)}),$
-which will turn out to be an equivalence by Theorem 4.9.
+$\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma);\quad M_{0}\mapsto(M_{0}\otimes_{A}A(\!(t^{\Gamma})\!),\operatorname{id}_{M_{0}\otimes_{A}A(\!(t^{\Gamma},u^{\Gamma})\!)}).$
+Remark 2.10.
+Our main theorem, Theorem 1.1, may now be reformulated to state that the functor $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma)$ is an equivalence of categories.
 Definition 2.11.
@@
 $x=\sum_{i\in I}a_{i}t_{1}^{d_{i,1}}\dotsm t_{n}^{d_{i,n}}$
-where $a_{i}\neq 0$ for all $i\in I$ and $(d_{i,1},\dotsc,d_{i,n})\in\Gamma^{n}$ are all distinct, we define the support of $x$ as
+where $a_{i}\neq 0$ for all $i\in I$ and $(d_{i,1},\dotsc,d_{i,n})\neq(d_{j,1},\dotsc,d_{j,n})$ for $i\neq j$ in $I$, we define the support of $x$ as
 $\operatorname{supp}(x)=\{(d_{i,1},\dotsc,d_{i,n}):i\in I\}\subseteq\Gamma^{n}.$
@@
 The support of $0\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ is empty.
-We always have $\operatorname{supp}(x+y)\subseteq\operatorname{supp}(x)\cup\operatorname{supp}(y)$ and $\operatorname{supp}(xy)\subseteq\operatorname{supp}(x)+\operatorname{supp}(y)$, where $+$ is the Minkowski sum defined as $X+Y=\{x+y:x\in X,y\in Y\}$.
+We always have $\operatorname{supp}(x+y)\subseteq\operatorname{supp}(x)\cup\operatorname{supp}(y)$ and $\operatorname{supp}(xy)\subseteq\operatorname{supp}(x)+\operatorname{supp}(y)$, where ++ is the Minkowski sum defined as $X+Y=\{x+y:x\in X,y\in Y\}$.
 By definition, for every element $x\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ or $x\in M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$, the intersection of $\operatorname{supp}(x)$ with the half-space
@@
 $f\colon M_{0}(\!(t^{\Gamma})\!)\to N_{0}(\!(t^{\Gamma})\!)$
-be a $A(\!(t^{\Gamma})\!)$-linear map compatible with the descent data on both sides.
+be an $A(\!(t^{\Gamma})\!)$-linear map compatible with the descent data on both sides. This means that the diagram
+$M_{0}(\!(t^{\Gamma},u^{\Gamma})\!)$$N_{0}(\!(t^{\Gamma},u^{\Gamma})\!)$$M_{0}(\!(t^{\Gamma},u^{\Gamma})\!)$$N_{0}(\!(t^{\Gamma},u^{\Gamma})\!)$$\scriptstyle{\pi_{1}^{\ast}f}$$\scriptstyle{\pi_{2}^{\ast}f}$
+commutes, where both $\pi_{1}^{\ast}f$ and $\pi_{2}^{\ast}f$ are $A(\!(t^{\Gamma},u^{\Gamma})\!)$-linear.
 Viewing $f$ as an element of $\operatorname{Hom}_{A}(M_{0},N_{0})(\!(t^{\Gamma})\!)$, we see that compatibility with descent data translates to
@@
 splits and hence is open by Lemma 2.14. This implies that there exists a large enough $N$ for which
-$t^{N}M(\!(t^{\mathbb{R}})\!)\subseteq\sum_{i=1}^{n}x_{i}A[[t^{\mathbb{R}}]].$
-In particular, for every $m\in M$, the element $t^{N}m\in M(\!(t^{\mathbb{R}})\!)$ can be written as $A[[t^{\mathbb{R}}]]$-linear combination of $x_{1},\dotsc,x_{n}$.
+$t^{N}M[[t^{\mathbb{R}}]]\subseteq\sum_{i=1}^{n}x_{i}A[[t^{\mathbb{R}}]].$
+In particular, for every $m\in M$, the element $t^{N}m\in M[[t^{\mathbb{R}}]]$ can be written as $A[[t^{\mathbb{R}}]]$-linear combination of $x_{1},\dotsc,x_{n}$.
 Therefore once we write $x_{i}=\sum_{j}m_{ij}t^{d_{ij}}$, the collection of $m_{ij}$ with $d_{ij}\leq N$ generate $M$, and so $M$ is finitely generated.
 We now choose an $A$-linear surjection
-$0\to K\to A^{\oplus m}\to M\to 0$
+$0\to K\to A^{\oplus m}\to M\to 0.$
 It follows that
@@
 A Novikov isocrystal over a ring $A$ is a pair $(M,F_{M})$, where $M$ is a finite projective $A(\!(t^{\mathbb{R}})\!)$-module and $F_{M}\colon M\to M$ is a $F$-semilinear bijection.
-A morphism $f\colon(M,F_{M})\to(N,F_{N})$ between two Novikov isocrystals is a $A(\!(t^{\mathbb{R}})\!)$-linear map $f\colon M\to N$ satisfying $f\circ F_{M}=F_{N}\circ f$.
+A morphism $f\colon(M,F_{M})\to(N,F_{N})$ between two Novikov isocrystals is an $A(\!(t^{\mathbb{R}})\!)$-linear map $f\colon M\to N$ satisfying $f\circ F_{M}=F_{N}\circ f$.
 We denote by $\mathsf{NovIsoc}(A)$ the category of Novikov isocrystals over $A$.
@@
 Proof.
-Let $M_{0}$ and $N_{0}$ be two finite projective $A$-modules, and let $f\colon M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)\to N_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)$ be a $A(\!(t^{\mathbb{R}})\!)$-linear map satisfying $f\circ(\operatorname{id}_{M_{0}}\otimes F)=(\operatorname{id}_{N_{0}}\otimes F)\circ f$.
+Let $M_{0}$ and $N_{0}$ be two finite projective $A$-modules, and let $f\colon M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)\to N_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)$ be an $A(\!(t^{\mathbb{R}})\!)$-linear map satisfying $f\circ(\operatorname{id}_{M_{0}}\otimes F)=(\operatorname{id}_{N_{0}}\otimes F)\circ f$.
 Regarding $f$ as an element of
@@
 we see that $f$ must be fixed under the automorphism $\operatorname{id}_{\operatorname{Hom}_{A}(M_{0},N_{0})}\otimes F$, and hence $\operatorname{supp}(f)=\lambda\operatorname{supp}(f)$.
-On the other hand, the intersection $[-1,1]\cap\operatorname{supp}(f)$ is finite, and therefore $\operatorname{supp}(f)\subseteq\{0\}$.
+On the other hand, the intersection $1,1]\cap\operatorname{supp}(f)$ is finite, and therefore $\operatorname{supp}(f)\subseteq\{0\}$.
 This shows that $f\in\operatorname{Hom}_{A}(M_{0},N_{0})$.
@@
 Because $k(\!(t^{\mathbb{R}})\!)$ is a field, it suffices to prove that for every nonzero Novikov isocrystal $(M,F_{M})$, there exists a nonzero element $m\in M$ and $c\in k^{\times}$ satisfying $F_{M}(m)=cm$.
-Fix an arbitrary nonzero $m\in M$ and find $n\geq 1$ such that $m,F_{M}(m),\dotsc,F_{M}^{n-1}(m)$ are linearly independent while
+Fix an arbitrary nonzero $m\in M$ and find $n\geq 1$ such that $m,F_{M}(m),\dotsc,F_{M}^{n-1}(m)$ are linearly independent over $k(\!(t^{\mathbb{R}})\!)$ while
 $F_{M}^{n}(m)=a_{0}m+a_{1}F_{M}(m)+\dotsb+a_{n-1}F_{M}^{n-1}(m)$
-for $a_{1},\dotsc,a_{n}\in k(\!(t^{\mathbb{R}})\!)$.
+for $a_{0},\dotsc,a_{n-1}\in k(\!(t^{\mathbb{R}})\!)$.
 Replacing $m$ with $t^{r}m$ for a real number $r\in\mathbb{R}$ has the effect of replacing $a_{i}$ with $t^{r(\lambda^{n}-\lambda^{i})}$.
-Thus by choosing $r$ appropriately, we may further assume that $a_{1},\dotsc,a_{n}\in k[[t^{\mathbb{R}}]]$ with $a_{i}\in k[[t^{\mathbb{R}}]]^{\times}$ for some $10\leq i\leq n-$ 1.
+Thus by choosing $r$ appropriately, we may further assume that $a_{1},\dotsc,a_{n}\in k[[t^{\mathbb{R}}]]$ with $a_{i}\in k[[t^{\mathbb{R}}]]^{\times}$ for some $0\leq i\leq n-1$.
 Our goal is to find elements $b_{0},\dotsc,b_{n-1}\in k(\!(t^{\mathbb{R}})\!)$, not all zero, and a scalar $c\in k^{\times}$ satisfying the equation
@@
 We claim that this equation has a nonzero solution in $k[[t^{\mathbb{R}}]]$.
-Let $\bar{a}_{i}\in k$ be the constant terms of $a_{i}\in k[[t^{\mathbb{R}}]]$.
-Since we are assuming that $\bar{a}_{i}\neq 0$ for some $11\leq i\leq n-$ 1 and that $k$ is algebraically closed, there exists a $c\in k^{\times}$ for which
+For each $0\leq i\leq n-1$, let $\bar{a}_{i}\in k$ be the constant term of $a_{i}\in k[[t^{\mathbb{R}}]]$.
+Since we are assuming that $\bar{a}_{i}\neq 0$ for some $1\leq i\leq n-1$ and that $k$ is algebraically closed, there exists a $c\in k^{\times}$ for which
 $\bar{a}_{0}+c\bar{a}_{1}+\dotsb+c^{n-1}\bar{a}_{n-1}=c^{n}.$
 Write
-$g(b)=c^{-n}F^{n}(b)F^{n-1}(a_{0})+c^{-n+1}F^{n-1}(b)F^{n-2}(a_{1})+\dotsb+c^{-1}F(b)$
+$g(b)=c^{-n}F^{n}(b)F^{n-1}(a_{0})+c^{-n+1}F^{n-1}(b)F^{n-2}(a_{1})+\dotsb+c^{-1}F(b)a_{n-1}$
 so that we need to find $b\in k[[t^{\mathbb{R}}]]$ with $g(b)=b$.
@@
 On the other hand, we observe that $x-y\in t^{d}k[[t^{\mathbb{R}}]]$ implies $g(x)-g(y)\in t^{\lambda d}k[[t^{\mathbb{R}}]]$.
-This shows that $b^{(j+1)}-b^{(j)}\in t^{\lambda^{i}\epsilon}k[[t^{\mathbb{R}}]]$ and therefore
+This shows that $b^{(j+1)}-b^{(j)}\in t^{\lambda^{j}\epsilon}k[[t^{\mathbb{R}}]]$ and therefore
 $b=\lim_{i\to\infty}b^{(j)}\in 1+t^{\epsilon}k[[t^{\mathbb{R}}]]$
@@
 $M\hookrightarrow M\otimes_{A(\!(t^{\mathbb{R}})\!)}B(\!(t^{\mathbb{R}})\!)\cong N_{0}\otimes_{B}B(\!(t^{\mathbb{R}})\!)=N_{0}(\!(t^{\mathbb{R}})\!)$
-to regard $M$ as a $A(\!(t^{\mathbb{R}})\!)$-submodule of $N_{0}(\!(t^{\mathbb{R}})\!)$ that is stable under the action of $\operatorname{id}_{N_{0}}\otimes F$.
+to regard $M$ as an $A(\!(t^{\mathbb{R}})\!)$-submodule of $N_{0}(\!(t^{\mathbb{R}})\!)$ that is stable under the action of $\operatorname{id}_{N_{0}}\otimes F$.
 Note that $A(\!(t^{\mathbb{R}})\!)\hookrightarrow B(\!(t^{\mathbb{R}})\!)$ is a closed subring with the subspace topology.
@@
 This is a closed $A(\!(t^{\mathbb{R}})\!)$-submodule of $(A^{\oplus n})(\!(t^{\mathbb{R}})\!)$ that is stable under $F^{\oplus n}$.
-Lemma 3.6 implies that there exists a $A$-submodule $K_{0}\subseteq A^{\oplus n}$ for which $K=K_{0}(\!(t^{\mathbb{R}})\!)$.
+Lemma 3.6 implies that there exists an $A$-submodule $K_{0}\subseteq A^{\oplus n}$ for which $K=K_{0}(\!(t^{\mathbb{R}})\!)$.
 On the other hand, $K$ has the subspace topology from $A(\!(t^{\mathbb{R}})\!)^{\oplus n}$, and therefore the $t$-adic topology on $K_{0}(\!(t^{\mathbb{R}})\!)$ agrees with the canonical topology.
@@
 In this subsection, we construct a functor $\mathsf{NovDD}(A,\mathbb{R})\to\mathsf{NovIsoc}(A)$. First note that for every finite projective module $M$ over $A(\!(t^{\mathbb{R}})\!)$, the diagram
-[…]
+$M{\pi_{2}^{\ast}M}$$\rho_{3}^{\ast}M$$\scriptstyle{\pi_{2}^{\ast}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}}$
 is an equalizer. Indeed, this can be checked for $M=A(\!(t^{\mathbb{R}})\!)$ by flatness of $M$, in which case it is clear.
@@
 For each $(M,\varphi)\in\mathsf{NovDD}(A)$, we define a $F_{2}$-semilinear bijection
-[…]
+$F_{M,2}\colon\pi_{2}^{\ast}M\xrightarrow{\varphi^{-1}}\pi_{1}^{\ast}M\xrightarrow{\operatorname{id}_{M}\otimes F_{2}}\pi_{1}^{\ast}M\xrightarrow{\varphi}\pi_{2}^{\ast}M.$
 Similarly, we can define a $F_{3}$-semilinear bijection
@@
 thanks to the cocycle condition. They fit in a commutative diagram
-[…]
-Both rows are equalizers, and all horizontal maps are $A(\!(t^{\mathbb{R}})\!)$-linear while the vertical maps are $F$-semilinear, if we use the ring homomorphisms $\pi_{2}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ and $\rho_{3}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$ to give every module a $A(\!(t^{\mathbb{R}})\!)$-module structure.
+$M{\pi_{2}^{\ast}M}$$\rho_{3}^{\ast}M$$M{\pi_{2}^{\ast}M}$$\rho_{3}^{\ast}M.$$\scriptstyle{\pi_{2}^{\ast}}$$\scriptstyle{F_{M}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}}$$\scriptstyle{F_{M,2}}$$\scriptstyle{F_{M,3}}$$\scriptstyle{\pi_{2}^{\ast}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}}$
+Both rows are equalizers, and all horizontal maps are $A(\!(t^{\mathbb{R}})\!)$-linear while the vertical maps are $F$-semilinear, if we use the ring homomorphisms $\pi_{2}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ and $\rho_{3}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$ to give every module an $A(\!(t^{\mathbb{R}})\!)$-module structure.
 Therefore we obtain an $F$-semilinear bijection $F_{M}\colon M\to M$, and this defines the functor
@@
 Let $A$ be any ring, and assume that the descent datum $(M,\varphi)\in\mathsf{NovDD}(A,\mathbb{R})$ induces $F_{M}\colon M\to M$. Then the diagram
-[…]
+$M{\pi_{2}^{\ast}M}$$\pi_{2}^{\ast}M$$\pi_{2}^{\ast}M$$\rho_{3}^{\ast}M$$\rho_{3}^{\ast}M$$\scriptstyle{\varphi\circ\pi_{1}^{\ast}}$$\scriptstyle{\varphi\circ\pi_{1}^{\ast}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{F_{M}\otimes F_{2}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}\varphi\circ\pi_{12}^{\ast}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{F_{M}\otimes F_{3}}$
 commutes, where both rows are equalizers.
@@
 Using the definition of $F_{M,2}$, we see that the first row being an equalizer is equivalent to
-[…]
+$M{\pi_{1}^{\ast}M}$$\pi_{1}^{\ast}M$$\scriptstyle{\pi_{1}^{\ast}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{\operatorname{id}_{M}\otimes F_{2}}$
 being an equalizer.
 Since all maps are $A(\!(t^{\mathbb{R}})\!)$-linear if we use $\pi_{1}^{\ast}$ to give $\pi_{1}^{\ast}M$ the structure of an $A(\!(t^{\mathbb{R}})\!)$-module, we may use finite projectivity of $M$ to reduce to
-[…]
+$A(\!(t^{\mathbb{R}})\!)$$A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$\scriptstyle{\pi_{1}^{\ast}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{F_{2}}$
 being an equalizer, which is now clear. The second row can similarly be seen to be an equalizer. ∎
@@
 Since $f\circ F_{M}=F_{N}\circ f$ and the right squares on the diagram only involve $F_{M},F_{N}$, the universal property of equalizers produces unique maps $g\colon M\to N$ and $h\colon\pi_{2}^{\ast}M\to\pi_{2}^{\ast}N$ fitting in the commutative diagram
-[…]
+$M{\pi_{2}^{\ast}M}$$N{\pi_{2}^{\ast}N}$$\pi_{2}^{\ast}M$$\rho_{3}^{\ast}M$$\pi_{2}^{\ast}N$$\rho_{3}^{\ast}N.$$\scriptstyle{\varphi_{M}\circ\pi_{1}^{\ast}}$$\scriptstyle{\varphi_{M}\circ\pi_{1}^{\ast}}$$\scriptstyle{g}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{2}^{\ast}f}$$\scriptstyle{\varphi_{N}\circ\pi_{1}^{\ast}}$$\scriptstyle{\varphi_{N}\circ\pi_{1}^{\ast}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}\varphi_{M}\circ\pi_{12}^{\ast}}$$\scriptstyle{h}$$\scriptstyle{\rho_{3}^{\ast}f}$$\scriptstyle{\pi_{23}^{\ast}\varphi_{N}\circ\pi_{12}^{\ast}}$
 We note that $g$ must be $A(\!(t^{\mathbb{R}})\!)$-linear, since all maps in the top face are $A(\!(t^{\mathbb{R}})\!)$-linear when we use $\pi_{1}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ for the $A(\!(t^{\mathbb{R}})\!)$-module structure.
@@
 $0=M_{0}\hookrightarrow M_{1}\hookrightarrow\dotsb\hookrightarrow M_{n}=M$
-with the proper that each associated graded $M_{i}/M_{i-1}$ is isomorphic to $(k(\!(t^{\mathbb{R}})\!),c_{i}F)$ for some $c_{i}\in k^{\times}$.
+with the property that each associated graded $M_{i}/M_{i-1}$ is isomorphic to $(k(\!(t^{\mathbb{R}})\!),c_{i}F)$ for some $c_{i}\in k^{\times}$.
 We inductively prove that each $(M_{i},F_{M}|_{M_{i}})\in\mathsf{NovIsoc}(k)$ is in the essential image of $\mathsf{Vect}(k)\to\mathsf{NovIsoc}(k)$, where the base case $i=0$ is clear.
@@
 has an $i$-dimensional kernel.
-Since the kernel when restricted to $\pi_{2}^{\ast}M_{i-1}$ is $(i-1)$-dimensional, there exist $x_{1},\dotsc,x_{i}\in k(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ with $x_{i}\neq 0$ satisfying
+Since the kernel when restricted to $\pi_{2}^{\ast}M_{i-1}$ is (i1)(i-1)-dimensional, there exist $x_{1},\dotsc,x_{i}\in k(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ with $x_{i}\neq 0$ satisfying
 $x_{1}=F_{2}(x_{1})+(\pi_{2}^{\ast}c_{1})F_{2}(x_{i}),\;\dotsc,\;x_{i-1}=F_{2}(x_{i-1})+(\pi_{2}^{\ast}c_{i-1})F_{2}(x_{i}),\;x_{i}=c_{i}F_{2}(x_{i}).$
@@
 and taking the product over all $i$ gives a commutative diagram
-[…]
+$\pi_{1}^{\ast}M\otimes_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$\pi_{2}^{\ast}M\otimes_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!).$$\scriptstyle{(\prod_{i}\pi_{1})^{\ast}f}$$\scriptstyle{\varphi\otimes\operatorname{id}}$$\scriptstyle{(\prod_{i}\pi_{2})^{\ast}f}$
 In particular, if we consider $M$ as an $A(\!(t^{\mathbb{R}})\!)$-submodule of $M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ via $f$, the images of
@@
 agree.
-For every element $\alpha=(\alpha_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, define the support of $\alpha$ as $\operatorname{supp}(\alpha)=\bigcup_{i\in I}\operatorname{supp}(\alpha_{i})\subseteq\mathbb{R}$, where the support of $\alpha_{i}\in M_{0,i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ is defined as before.
-For $\beta=(\beta_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ we similarly define $\operatorname{supp}(\beta)=\bigcup_{i}\operatorname{supp}(\beta_{i})\subseteq\mathbb{R}$.
+For every element $\alpha=(\alpha_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, define the support of $\alpha$ as $\operatorname{supp}(\alpha)=\bigcup_{i\in I}\operatorname{supp}(\alpha_{i})\subseteq\mathbb{R}^{2}$, where the support of $\alpha_{i}\in M_{0,i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ is as in Definition 2.11.
+For $\beta=(\beta_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ we similarly define $\operatorname{supp}(\beta)=\bigcup_{i}\operatorname{supp}(\beta_{i})\subseteq\mathbb{R}^{2}$.
 We now claim that for each $m\in M\subseteq M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$, the support $\operatorname{supp}(m)\subseteq\mathbb{R}$ is bounded below.
@@
 $\operatorname{supp}(m_{j,i})\times\{0\}\subseteq\bigcup_{k=1}^{n}(\operatorname{supp}(a_{jk,i})+(\{0\}\times\operatorname{supp}(m_{k,i})))$
-inside $\mathbb{R}^{2}$, where by $+$ we mean the Minkowski sum.
+inside $\mathbb{R}^{2}$, where by ++ we mean the Minkowski sum.
 By our assumptions, the right hand side is always contained in $\{(x,y):2x+y\geq C+D_{i}\}$.
@@
 $\ker(\mathrm{GL}_{A}(M_{i}|_{U_{i}})\to\mathrm{GL}_{A/I}(\bar{M}|_{U_{i}}))\cong\operatorname{Hom}_{A/I}(\bar{M},I\otimes_{A/I}\bar{M})|_{U_{i}}.$
-Since the Čech $H^{2}$ of quasi-coherent sheaves with respect to $\{U_{i}\}$ vanish on $\operatorname{Spec}A/I$ as each $U_{i}$ is affine, this is a coboundary, and hence we may modify the isomorphisms $M_{i}|_{U_{i}\cap U_{j}}\cong M_{j}|_{U_{i}\cap U_{j}}$ so that they do satisfy the cocycle condition.
+The second Čech cohomology group of the quasi-coherent sheaf $\operatorname{Hom}_{A/I}(\bar{M},I\otimes_{A/I}\bar{M})$ with respect to $\{U_{i}\}$ vanishes on $\operatorname{Spec}A/I$ as each $U_{i}$ is affine.
+Therefore the $2$-cocycle is a coboundary, and hence we may modify the isomorphisms $M_{i}|_{U_{i}\cap U_{j}}\cong M_{j}|_{U_{i}\cap U_{j}}$ so that they do satisfy the cocycle condition.
 ∎
@@
 $\operatorname{Hom}_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)}(M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!),(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!))$ $\cong\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!),$
-$\operatorname{Hom}_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)}(M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!),(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!))$ $\cong\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!).$
+$\operatorname{Hom}_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)}(M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!),(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!))$ $\cong\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$
+by Lemma 2.8.
 Using this isomorphism, we identify $\psi_{+}$ as an element of $\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, and similarly the additive cocycle condition as an equation valued in the group $\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$.
-By considering the coefficients in the cocycle condition, we readily see that $\psi_{+}$ is supported on ({0})×{0}{0}×($(\mathbb{R}\setminus\{0\})\times\{0\}\cup\{0\}\times(\mathbb{R}\setminus\{0\})$, and moreover the coefficient of $t^{d}$ agrees with the coefficient of $u^{d}$ for each $d\in\mathbb{R}\setminus\{0\}$.
-This allows us to write
+We now note that the cocycle condition implies
+$\{(x,0,z):(x,z)$ $\in\operatorname{supp}(\psi_{+})\}=\operatorname{supp}(\pi_{13}^{\ast}\psi_{+})$
+$\subseteq\operatorname{supp}(\pi_{23}^{\ast}\psi_{+})\cup\operatorname{supp}(\pi_{12}^{\ast}\psi_{+})\subseteq(\{0\}\times\mathbb{R}^{2})\cup(\mathbb{R}^{2}\times\{0\}).$
+It follows $\operatorname{supp}(\psi_{+})$ is contained in $(\{0\}\times\mathbb{R})\cup(\mathbb{R}\times\{0\})$. This allows us to write
+$\psi_{+}=c+\sum_{0\neq d\in\mathbb{R}}(a_{d}t^{a}+b_{d}u^{d}),\quad c,a_{d},b_{d}\in\operatorname{Hom}_{A}(M_{0},IM_{0}).$
+The cocycle condition then can be written out as
+$c+\sum_{0\neq d\in\mathbb{R}}(a_{d}t^{d}+b_{d}v^{d})=2c+\sum_{0\neq d\in\mathbb{R}}(a_{d}t^{d}+(b_{d}+a_{d})u^{d}+b_{d}v^{d})$
+and hence $c=0$ and $b_{d}=-a_{d}$ for all $0\neq d\in\mathbb{R}$. Setting $\xi_{+}=\sum_{0\neq d\in\mathbb{R}}a_{d}t^{d}$, we obtain
 $\psi_{+}=\pi_{1}^{\ast}\xi_{+}-\pi_{2}^{\ast}\xi_{+},\quad\xi_{+}\in\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}})\!).$
-Regarding $\xi_{+}$ as a $A(\!(t^{\mathbb{R}})\!)$-linear map $M_{0}(\!(t^{\mathbb{R}})\!)\to(IM_{0})(\!(t^{\mathbb{R}})\!)$, we now define $\xi=\operatorname{id}+\xi_{+}$. It follows that
+Regarding $\xi_{+}$ as an $A(\!(t^{\mathbb{R}})\!)$-linear map $M_{0}(\!(t^{\mathbb{R}})\!)\to(IM_{0})(\!(t^{\mathbb{R}})\!)$, we now define $\xi=\operatorname{id}+\xi_{+}$. It follows that
 $\psi=\operatorname{id}+\pi_{1}^{\ast}\xi_{+}-\pi_{2}^{\ast}\xi_{+}=(\operatorname{id}-\pi_{2}^{\ast}\xi_{+})\circ(\operatorname{id}+\pi_{1}^{\ast}\xi_{+})=\pi_{2}^{\ast}\xi^{-1}\circ\pi_{1}^{\ast}\xi.$
@@
 $x=\sum_{i=1}^{n}(\pi_{1}^{\ast}m_{i})a_{i},\quad m_{i}\in M,\quad a_{i}\in A(\!(t^{\Gamma},u^{\Gamma})\!),$
-we see that $\operatorname{supp}(x)$ must be contained in ×$\mathbb{R}\times\Gamma$.
+we see that $\operatorname{supp}(x)$ must be contained in $\mathbb{R}\times\Gamma$.
 For every element $m\in M$, we in particular have $\operatorname{supp}(\pi_{2}^{\ast}m)\subseteq\mathbb{R}\times\Gamma$, and hence $\operatorname{supp}(m)\subseteq\Gamma$.
@@
 This shows that the composition
-[…]
+$M_{0}(\!(t^{\Gamma})\!)\xrightarrow{h^{\vee}}M\xrightarrow{g}M_{0}(\!(t^{\Gamma})\!)$
 is the identity map upon base changing along $A(\!(t^{\Gamma})\!)\to{A(\!(t^{\mathbb{R}})\!)}$.
@@
 For a v-sheaf $X$, we define a vector bundle on $X$ to be an association to each $(R,R^{+})$-point $\xi\in X(R,R^{+})$ a locally finite free sheaf $\mathscr{V}_{\xi}$ on $\operatorname{Spa}(R,R^{+})$, together with the data of, for each map $f\colon\operatorname{Spa}(S,S^{+})\to\operatorname{Spa}(R,R^{+})$ and $\xi\in X(R,R^{+})$, an isomorphism $f^{\ast}\mathscr{V}_{\xi}\cong\mathscr{V}_{f^{\ast}\xi}$, satisfying the condition that for every
-[…]
+$\operatorname{Spa}(T,T^{+})\xrightarrow{g}\operatorname{Spa}(S,S^{+})\xrightarrow{f}\operatorname{Spa}(R,R^{+})$
 and $\xi\in X(R,R^{+})$, the three identifications between $g^{\ast}f^{\ast}\mathscr{V}_{\xi}$, $g^{\ast}\mathscr{V}_{f^{\ast}\xi}$, and $\mathscr{V}_{g^{\ast}f^{\ast}\xi}$ commute.
+Let $A$ be a discrete perfect $\mathbb{F}_{p}$-algebra. There is a v-cover
+$X=\operatorname{Spd}(A(\!(t^{{1/p^{\infty}}})\!),A[[t^{{1/p^{\infty}}}]])\to\operatorname{Spd}(A,A)=Y$
+of $Y$ be an affinoid perfectoid space $X$.
+We note that $X$ represents the functor sending a perfectoid Huber pair $(A,A)\to(R,R^{+})$ to a topologically nilpotent unit $\varpi\in R^{\circ\circ}$, because uniformity and perfectness of $R$ implies that there exists a unique continuous map $A(\!(t^{{1/p^{\infty}}})\!)\to R$ sending $t\mapsto\varpi$.
+This means that both $X\times_{Y}X$ and $X\times_{Y}X\times_{Y}X$ are representable by perfectoid spaces
+$X\times_{Y}X$ $=\operatorname{Spa}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\setminus\{\lvert tu\rvert=0\}$
+$=\bigcup_{n=2}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl\langle\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr\rangle[1/tu],$
+$X\times_{Y}X\times_{Y}X$ $=\operatorname{Spa}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}}]]\setminus\{\lvert tuv\rvert=0\}$
+$=\bigcup_{n=3}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}}]]\Bigl\langle\frac{t^{n}}{uv},\frac{u^{n}}{tv},\frac{v^{n}}{tu}\Bigr\rangle[1/tuv].$
+Lemma 5.2.
+We have
+$H^{0}(X\times_{Y}X,\mathscr{O}_{X\times_{Y}X})$ $=A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!),$
+$H^{0}(X\times_{Y}X\times_{Y}X,\mathscr{O}_{X\times_{Y}X\times_{Y}X})$ $=A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}})\!).$
+Proof.
+For the first equality, we note that we can write
+$A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl\langle\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr\rangle\Bigl[\frac{1}{tu}\Bigr]=\left\{\sum_{i\in I}a_{i}t^{d_{i}}u^{e_{i}}:\begin{matrix}a_{i}\in A,d_{i},e_{i}\in\mathbb{Z}[p^{-1}],\\ \{i\in I:\alpha d_{i}+e_{i}<C\}\text{ is finite}\\ \text{for all }C\in\mathbb{R}\text{ and }n^{-1}\leq\alpha\leq n\end{matrix}\right\}.$
+As we increase the $n$ and take the intersection of the rings, we see that we allow all possible values of $\alpha>0$, and hence the ring of global sections of $X\times_{Y}X$ is $A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!)$.
+The proof of the second equality is similar.
+∎
 Corollary 5.3.
-Let $A$ be a perfect $\mathbb{F}_{p}$-algebra. Then the natural functor
-{vector bundles on $\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
+Let $A$ be a discrete perfect $\mathbb{F}_{p}$-algebra. Then the natural functor
+$\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
 is an exact tensor equivalence.
 Proof.
-As the natural map
+We use the v-cover
 $X=\operatorname{Spd}(A(\!(t^{{1/p^{\infty}}})\!),A[[t^{{1/p^{\infty}}}]])\to\operatorname{Spd}A=Y$
-is a v-cover, by [SW20, Lemma 17.1.8] the category of vector bundles on $\operatorname{Spd}A$ can be identified with the category of pairs $(\mathscr{V},\varphi)$, where $\mathscr{V}$ is a vector bundle on $X$ and $\varphi$ is an isomorphism of vector bundles over $X\times_{Y}X$ satisfying a cocycle condition over $X\times_{Y}X\times_{Y}X$.
+from above.
+By [SW20, Lemma 17.1.8] the category of vector bundles on $\operatorname{Spd}A$ can be identified with the category of pairs $(\mathscr{V},\varphi)$, where $\mathscr{V}$ is a vector bundle on $X$ and $\varphi$ is an isomorphism of vector bundles over $X\times_{Y}X$ satisfying a cocycle condition over $X\times_{Y}X\times_{Y}X$.
 By [KL15, Theorem 2.7.7] and again [SW20, Lemma 17.1.8], the vector bundle $\mathscr{V}$ corresponds to a finite projective module $M$ over $A(\!(t^{{1/p^{\infty}}})\!)$.
-Next, we identify
-$X\times_{Y}X=\bigcup_{n=2}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr{\rangle}[1/tu].$
-If we denote by $\pi_{1},\pi_{2}\colon X\times_{Y}X\to X$ the two projection maps, then $\varphi$ is an isomorphism of vector bundles
-$\varphi\colon\pi_{1}^{\ast}\mathscr{V}\xrightarrow{\cong}\pi_{2}^{\ast}\mathscr{V}$
-over $X\times_{Y}X$. Because
-$H^{0}(X\times_{Y}X,\mathscr{O}_{X\times_{Y}X})=\bigcap_{n=2}^{\infty}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr{\rangle}[1/tu]=A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!),$
-we see that $\varphi$ corresponds to an isomorphism
+By Lemma 5.2, the natural functor from the category of finite projective $A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!)$-modules to the category of vector bundles on $X\times_{Y}X$ is fully faithful.
+If we denote by $\pi_{1},\pi_{2}\colon X\times_{Y}X\to X$ the two projection maps, then the descent datum $\varphi\colon\pi_{1}^{\ast}\mathscr{V}\cong\pi_{2}^{\ast}\mathscr{V}$ corresponds to an isomorphism of $A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!)$-modules
 $\varphi\colon M\otimes_{A(\!(t^{{1/p^{\infty}}})\!),\pi_{1}^{\ast}}A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!)\xrightarrow{\cong}M\otimes_{A(\!(t^{{1/p^{\infty}}})\!),\pi_{2}^{\ast}}A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!).$
-Finally the cocycle condition is defined on
-$X\times_{Y}X\times_{Y}X=\bigcup_{n=3}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{uv},\frac{u^{n}}{tv},\frac{v^{n}}{tu}\Bigr{\rangle}[1/tuv],$
-and hence be checked after tensoring to
+Finally the cocycle condition is defined on $X\times_{Y}X\times_{Y}X$, and hence by Lemma 5.2 can be checked after tensoring to
 $H^{0}(X\times_{Y}X\times_{Y}X,\mathscr{O}_{X\times_{Y}X\times_{Y}X})=A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}})\!).$
-This shows that a vector bundle on $\operatorname{Spd}A$ corresponds to a Novikov descent datum for $A$. Applying Theorem 4.9 now gives the desired result. ∎
+This shows that a vector bundle on $\operatorname{Spd}A$ corresponds to a Novikov descent datum for $A$ with $\Gamma=\mathbb{Z}[p^{-1}]$. Applying Theorem 1.1 now gives the desired result. ∎
 We now compare the finite étale sites of $\operatorname{Spec}A$ and $\operatorname{Spd}A$.